### v1

# Building Confidence in Deep Generative Protein Design

Abstract
Deep generative models show promise for de novo protein design, but their effectiveness within specific protein families remains underexplored. In this study, we evaluate two 3D rigid-body generative methods, score matching and flow matching, to generate monomeric protein backbones in $\mathrm{SE}(3)$ space. Our goal is to provide new insights and build confidence in the broader applicability of deep generative models for protein design. The optimal amino acid sequences were predicted from the generated backbones, followed by side-chain homology modeling. Results demonstrated high structural integrity, with conserved key residues aligning with known proteins. Structural phylogenetic analysis shows evolutionary links between the generated samples and their protein family members. Further molecular dynamics simulations and protein-ligand docking confirm the dynamic stability and functional potential of these samples, with ligand binding inducing conformational changes consistent with those in wild-type proteins.
Machine Learning
1 Introduction
Protein functions and specificity are determined by their complex structures. Over the past 60 years, we have progressed from viewing protein design as unattainable to achieving complete artificial design and synthesis (Korendovych & DeGrado, 2020; Huang et al., 2016), with expanding applications across industries (Arunachalam et al., 2021; Kingwell, 2024; Barclay & Acharya, 2023). However, precise design remains challenging due to the nonlinear complexity of protein folding and the sensitivity of function to slight changes, still demanding significant resources.
With the rapid growth of protein structure databases (Jamasb et al., 2024; Berman, 2000), various in silico de novo protein design methods, particularly deep generative models (Watson et al., 2023; Ingraham et al., 2023; Wu et al., 2022), have emerged. Approaches such as diffusion-based score matching (SM) (Yim et al., 2023), and flow matching (FM) (Bose et al., 2023), show competitive performance for generating monomeric protein backbones in $\mathrm{SE}(3)$ [^1] Special Euclidean group. See more in Section A.1. space.
While many methods show novelty, diversity, and designability in generated structures, few studies assess their functionality or evolutionary relevance, leaving the observed novelty and diversity potentially as mere hallucinations (Ji et al., 2023). Unclear biological functionality, unknown stability, and the lack of validation connecting these structures to known functions limit the broader application and advancement of generative methods.
This study evaluates the effectiveness of SM and FM methods across four well-studied protein families ($\beta$-lactamases, cytochrome c, GFP, and Ras). After determining optimal amino acid sequences and conducting side-chain homology modeling, we assess the functional relevance and evolutionary potential of the design, aiming to build confidence in deep generative approaches for protein design.
Our key contributions are: (1) We present a deep generative model pipeline for early-stage de novo protein design, with a robust evaluation protocol adaptable to a wider range of protein structure generation methods. (2) Despite sequence dissimilarity, generated samples show structural similarity to experimentally determined proteins in their families, conserving key residues linking to functions. (3) Structural phylogenetic analysis of both experimental and generated proteins reveals evolutionary relationships, showing the potential of structural phylogenetics. (4) Molecular dynamics (MD) simulations and protein-ligand docking show the dynamic stability and functional potential of generated samples. These samples form binding pockets akin to wild-type (WT) proteins, with conserved residues that bind family-specific ligands at low energies. For instance, generated KRas-like structures exhibit switch region flexibility and GTP/GDP-dependent conformational shifts similar to WT. (5) FM samples show greater flexibility, diversity, and novelty, while SM better captures conserved regions, offering insights for model selection and optimization in future work.
2 Background: Protein Backbone Generation
In this section and Appendix A, we review the key concepts behind Yim et al. (2023) and Bose et al. (2023)’s approaches for backbone generation, using SM and FM on $\mathrm{SE}(3)$.
Protein backbone representations.
Molecules can be intuitively represented as atomic point clouds; however, for macromolecules, this approach often leads to high dimensionality and data sparsity. A more compact, fragment-based representation involves using backbone rigid groups from AlphaFold (Jumper et al., 2021). For each amino acid residue $i\in[1,N]$, its group consists of the backbone atoms $[\mathrm{N},\mathrm{C}_{\alpha},\mathrm{C},\mathrm{O}]_{i}$ (Figure 1A). The goal is to learn how the rigid transformation (or frame) $\bm{\mathsf{T}}_{i}$ acts on the idealized coordinates, so that the transformed coordinates match the actual coordinates. For simplicity, $\bm{\mathsf{T}}_{i}\in\mathrm{SE}(3)$ decomposes into a rotation matrix $\bm{\mathsf{R}}_{i}\in\mathrm{SO}(3)$ and a translation vector $\bm{\mathsf{x}}_{i}\in\mathbb{R}^{3}$. An extra dihedral angle $\psi_{i}\in\mathrm{SO}(2)$ is introduced between the $\mathrm{C}_{\alpha}$ and C bond to optimize the positioning of the backbone oxygen ($\mathrm{O}$).
$\mathrm{SE}(3)$ score matching.
Let $\mathbf{T}_{t}=\left[\bm{\mathsf{T}}_{1,t},\dots,\bm{\mathsf{T}}_{N,t}\right]$ be the manifold of $N$ frames at time $t$. The forward process gradually perturbs the distribution $\mathbf{T}_{0}\sim p_{0}$, pushing it toward a simple distribution $p_{T}$, such as a Gaussian, by introducing a stochastic process independently in rotation and translation spaces (Equation 4). The reverse process approximately reconstructs $p_{0}$ by iteratively calculating the gradient of the log probability density (or the score) $\nabla\log p$ along a reverse path (Equation 5 and Equation 6). Since this score function is typically intractable in practice, a network $s(\theta,t,\cdot)$ is trained to approximate the conditional score $\nabla\log p_{t|0}$ at time $t\sim\mathcal{U}(0,T)$ (Vincent, 2011).
$\mathrm{SE}(3)$ flow matching.
The flow from $p_{T}$ to $p_{0}$ is defined as a set of continuously differentiable, time-indexed vector fields $\{\mathbf{u}_{t}\}$, where FM learns the time evolution of $\mathbf{u}_{t}$ to gradually transport the data from $p_{T}$ to $p_{0}$. Because $\mathbf{u}_{t}$ is generally intractable, Lipman et al. (2022) introduced conditional FM, training a neural network $v(\theta,t,\cdot)$ to approximate $\mathbf{u}_{t|0}$ (Equation 13). When the optimal transport condition is satisfied, it provides a smooth and efficient transformation path, improving training stability and reducing computational complexity (Pooladian et al., 2023).
3 Evaluations of Generated Structures
Figure 1: (A) Protein backbone and dihedral angles $\psi$ and $\phi$. (B) Ramachandran plots showing $\psi$ and $\phi$ angle distributions for generated protein backbones vs. experimental proteins. The top-right inset shows conformationally favored (light) and disallowed (dark) angle regions. (C) Structural overlay of WT E. coli TEM1 (PDB: 1BTL; Jelsch et al. (1993); green) and its E166N acylated intermediate (PDB: 1FQG; Brown et al. (2009); white) with penicillin (PNM). (D) Structure of the WT E. caballus heart cytochrome c (PDB: 1HRC; (Dickerson et al., 1967)) with heme C (HEC). (E) Structure of A. victoria GFP and its chromophore (PDB: 4KW4; Barnard et al. (2014)). (F) Structure of GDP-bound H. sapiens KRas protein (PDB: 4OBE; (Hunter et al., 2014)).
3.1 Data
We generated monomeric proteins across a range of families, covering diverse fold types and functions. These proteins include both natural and engineered mutations while preserving conserved core functional regions.
$\beta$-lactamases.
$\beta$-lactamases are enzymes that deactivate $\beta$-lactam antibiotics by hydrolyzing their $\beta$-lactam ring, contributing significantly to bacterial resistance (Lee et al., 2016). Using $\beta$-lactamase inhibitors can restore antibiotic efficacy by blocking this reaction (Behzadi et al., 2020). Since their discovery, $\beta$-lactamase diversity has rapidly expanded due to the evolution of bacterial resistance, making them ideal for protein modeling studies. For this study, we gathered structural data on 1,578 unique monomeric $\beta$-lactamases and their variants across Ambler classes (A, B, C, and D) from the BLDB (Naas et al., 2017) and the Protein Data Bank (PDB) (Berman, 2000). Class A $\beta$-lactamases are the most prevalent, with conserved active-site residues Ser70, Glu166, and Asn170 coordinating the hydrolytic water for deacylation (Tooke et al., 2019; Brown et al., 2009). Figure 1C shows the wild-type $\beta$-lactamase TEM1 (1BTL) alongside its acylated E166N intermediate (1FQG).
Cytochrome c.
Cytochrome c is a water-soluble protein ($\sim 12$ kDa) essential for ATP synthesis in mitochondria and intrinsic apoptosis (Kashyap et al., 2021; Ow et al., 2008). It also serves as an independent marker for apoptosis in several cancers (Li et al., 2001; Way et al., 2004). Despite variations across species, its core structure and function are conserved. We obtained structural data for 498 unique cytochrome c proteins and variants from the PDB.
Figure 1D shows horse cytochrome c, where a hydrophobic shell surrounds the heme group, with only $\sim 7.5\%$ of the surface available for electron transfer (Bushnell et al., 1990). The hydrophobic environment and iron coordination by His18 and Met80 maintain a high redox potential ( 260 mV) (Salemme, 1977). Phosphorylation occurs at Thr28, Thr47, Tyr48, and Tyr97 (Hüttemann et al., 2011), while Lys72, Lys73, and Lys87 bind phospholipids (Kagan et al., 2009). The ATP-binding pocket includes Glu69, Asn70, Lys88, and Lys72, Lys86, Lys87 (McIntosh et al., 1996).
Green fluorescent proteins (GFP).
GFP, first isolated from Aequorea victoria, emits green fluorescence under specific wavelengths. Its core structure is an 11-strand $\beta$-barrel that encloses the chromophore (Figure 1E) (Remington, 2011). Various mutants have been engineered to enhance or modify its properties, including enhanced GFP (Cormack et al., 1996), superfolder GFP (Pédelacq et al., 2005), and color variants like YFP (Ormö et al., 1996) and BFP (Glaser et al., 2016). We collected structural data for 448 GFPs and variants from the PDB to explore this diversity.
Ras.
Ras proteins, a subgroup of the small GTPase superfamily, act as molecular switches, cycling between GTP-bound (active) and GDP-bound (inactive) states to regulate cell proliferation, differentiation, migration, and apoptosis (Ladygina et al., 2011; Simanshu et al., 2017; Weinmann & Ottow, 2007). They play a key role in signaling from the cell surface to downstream pathways. Mutations that keep Ras proteins in an active state drive excessive cell growth and malignancy, making Ras inhibition a promising cancer treatment strategy (Singh & Lingham, 2002). We focused on the most common cancer-related Ras proteins (HRas, KRas, and NRas) (Cox, 2002) and obtained 511 experimental structures from the PDB.
In human KRas (Figure 1F), the switch I and II regions form the key interface for interacting with effectors and regulators (Pantsar, 2020). These regions are highly flexible, with conformations depending on GTP or GDP binding. Cancer-related mutations frequently occur in the P-loop and switch II (Pantsar, 2020).
3.2 Training and Sampling
We trained the SM and FM (with OT) models on the described protein families, using pretrained weights from Yim et al. (2023) and Bose et al. (2023), with each model having $\sim 17$ million parameters. The sequence length distributions of the training proteins are shown in Figure 6. Each model generated 50 backbone structures with target lengths of 270 amino acids for $\beta$-lactamases, 113 for cytochrome c, 238 for GFP, and 170 for Ras proteins.
3.3 Backbone Dihedral Angles
Due to steric hindrance and spatial repulsion, not all backbone dihedral angles in proteins are physically feasible or energetically favorable. We analyzed the $\phi$ and $\psi$ distribution in the generated structures using Ramachandran plots (Ramachandran et al., 1963) (Figure 1B). Most data points fall within the allowed and favored regions, with only a few in the disallowed areas, and no significant geometric clashes or unreasonable conformations were observed.
The dihedral angle distributions of the generated structures align well with those of the experimental training proteins. For instance, cytochrome c structures have sparse points in the region $-180^{\circ}<\psi<-90^{\circ}$ and $45^{\circ}<\phi<180^{\circ}$, while GFP structures form four clusters there. SM samples are closer to the training data and concentrate in allowed regions but show less diversity than FM samples.
3.4 Conserved Residue Consistency
Figure 2: (A) Normalized evolutionary rates mapped onto structures, with black for rapidly evolving positions and orange for conserved ones. Experimental structures are highlighted with bold outlines, and highly conserved residues are marked in red (exposed) and green (buried). (B) Summary tree of two structural phylogenetic trees (Figure 7), constructed using $Q_{\text{score}}$ and the 3Di alphabet. Protein family members are grouped in semi-circular rings, with specific subtypes labeled in white. In $beta$-lactamases, different Ambler classes are shown with colored backgrounds. Colored branches and nodes indicate generated protein structures.
In protein families or across species, certain residues are highly conserved, typically to ensure structural stability and support proper folding and function.
Determining the optimal amino acid sequence.
Following Yim et al. (2023), we used ProteinMPNN (Dauparas et al., 2022) to predict ten sequences for each generated backbone. We then modeled these sequences with EMSfold (Rives et al., 2019) and compared them to the original backbone using the $\text{TM}_{\text{score}}$ (Zhang, 2005). The sequence with the highest $\text{TM}_{\text{score}}$ was chosen as the optimal match.
Identifying conserved residues.
The experimental protein sequences were aligned with each generated backbone’s optimal sequence using Clustal Omega (Sievers et al., 2011). Given these multiple sequence alignments (MSAs) and generated structures, ConSurf (Yariv et al., 2023) reconstructed phylogenetic trees and applied Rate4Site (Pupko et al., 2002) to calculate evolutionary rates for each position with an empirical Bayesian method (Ashkenazy et al., 2016; Mayrose, 2004). The optimal predicted sequences for the generated backbones showed conserved residues matching those in the experimental sequences (Figure 8). FM samples displayed greater diversity than SM samples, suggesting higher novelty, as noted in Section 3.5. Interestingly, this trend reverses in generated $\beta$-lactamases, possibly due to the greater structural variability in the experimental $\beta$-lactamase training set (Figure 2B), to which SM models are more sensitive. Section 3.6 further assesses the dynamic stability and ligand-binding properties of these generated samples.
In Figure 2A, similar to the experimental proteins, the generated structures have higher residue conservation near binding pockets. For instance, in cytochrome c (1HRC) and generated structures SM-1 and FM-0, conserved residues cluster around the central heme C binding site. A similar pattern is seen in 4OBE and the generated KRas proteins.
3.5 Structural Phylogenetics
In specific tasks, we aim to generate structures with dissimilar sequences while preserving similar functions. However, when sequence similarity falls below 30%, detecting homology and evolutionary relationships becomes challenging (Puente-Lelievre et al., 2023). Structural comparison, being more conserved (Illergård et al., 2009), is more effective for uncovering evolutionary connections (Flores et al., 1993; Puente-Lelievre et al., 2023; Moi et al., 2023).
Using $Q_{\text{score}}$ in structural phylogenetics.
Malik et al. (2020) proposed the $Q_{\text{score}}$ (Krissinel & Henrick, 2004) for structural phylogenetics, as it accommodates indels and combines alignment quality with length. $Q_{\text{score}}$ compares the positions of all $\mathrm{C}_{\alpha}$ atoms among $N_{\text{align}}$ comparable residues in pairwise comparisons. We construct structural phylogenetic trees using $1-Q_{\text{score}}$ as a distance measure, with higher values indicating greater structural similarity. For any two structures with $N_{1}$ and $N_{2}$ residues, we align and calculate $Q_{\text{score}}$ using TM-align (Zhang, 2005) as:
$Q_{\text{score}}=\frac{N_{\text{align}}^{2}}{N_{1}N_{2}}\times\frac{1}{1+\left(\frac{\text{RMSD}}{R_{0}}\right)^{2}}$ (1)
[…]
Using 3Di alphabet in structural phylogenetics.
van Kempen et al. (2023) developed Foldseek, which represents protein tertiary interactions as sequences over a 20-state 3D interaction (3Di) alphabet, simplifying structural alignments. This 3Di alphabet reduces false positives and increases information density by encoding conserved core regions more effectively. Leveraging Foldseek’s structural divergence metrics, Moi et al. (2023) constructed structural phylogenetic trees based on rigid body alignment, local alignment without superposition, and sequence alignment with structural alphabets, showing that these trees outperform traditional sequence-based trees across varied evolutionary timescales (Moi et al., 2023).
Summary tree.
Two methods were used to calculate pairwise structural distances between generated and experimental structures, creating a distance matrix for constructing structural phylogenetic trees (Figure 7). After normalizing branch lengths, we generated a summary tree using SumTrees (Moreno et al., 2024), shown in Figure 2B. To ensure reliable $Q_{\text{score}}$ calculations, residue count differences between compared structures were limited to within 10% (Malik et al., 2020). Thus, we included proteins with residue counts of 243-297 for $\beta$-lactamases, 102-124 for cytochrome c, 216-261 for GFP, and 166-187 for Ras proteins, excluding class C $\beta$-lactamases ($\sim 360$ residues) from the tree but retaining them for model training and analysis.
The summary tree illustrates the potential of structural phylogenies, with each protein family clustering on distinct branches and subtypes (such as metallo-$\beta$-lactamases) grouping closely. Generated structures mostly integrate within small branches of experimental proteins, suggesting a close structural relationship. However, the models may not fully capture the specificity of class D $\beta$-lactamases, as no generated samples clustered in this class.
SM samples form dense clusters, while FM samples show greater diversity, positioning closer to the tree root but near experimental proteins without creating new branches, indicating higher novelty in FM outputs.
3.6 Molecular Dynamics
Figure 3: Stability assessment of MD simulations across proteins using various metrics. (A, C-E) Distributions of (A) RMSD, (C) radius of gyration ($R_{g}$), (D) potential energy, and (E) secondary structure counts throughout the simulation. Interquartile ranges and whiskers show metric variation; high-quality structures have medians close to experimental values with narrow ranges. (B) Residue fluctuation (RMSF) during simulation, where stable structures exhibit lower RMSF values and trends similar to the experimental structure.
Structures that appear reasonable may, in fact, be unstable due to molecular dynamics, water interactions, and entropy, which are not considered during generation. To evaluate the dynamic stability of the generated structures, we conducted MD simulations under physiological conditions and analyzed their time-dependent behavior.
Homology modeling of side-chains.
We used homology modeling to add side-chains to the protein backbone. After determining the optimal sequence (Section 3.4), we selected template proteins with at least 45% sequence identity and a $\text{TM}_{\text{score}}$ of 0.75 or higher using FoldSeek (van Kempen et al., 2023). Sequences were aligned with Clustal Omega, and MODELLER (Šali & Blundell, 1993) generated possible side-chain conformations using statistical potentials and rotamer libraries, with backbone fixed. We chose the final side-chain arrangement based on lowest energy and minimal steric clashes, verifying model quality with PROCHECK (Laskowski et al., 1993, 1996) and WHAT_CHECK (Hooft et al., 1996), discarding low-quality models.
Simulation.
After adding hydrogen atoms with Reduce2 from the computational crystallography toolbox (Grosse-Kunstleve et al., 2002), we conducted MD simulations with GROMACS (Abraham et al., 2015). The protein was placed in an octahedral box with at least 15 Å̊$\mathring{\mathrm{A}}$ from the edges, using the CHARMM36 force field (July 2022) (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Yu et al., 2012; Soteras Gutiérrez et al., 2016) for intermolecular interactions. Following vacuum energy minimization, the system was solvated, neutralized to 150 mM Na+ and Cl-, and minimized again. The system was then heated to 310 K under NVT [^2] Constant number of particles, volume, and temperature. conditions and equilibrated at 1 atm under NPT [^3] Constant number of particles, pressure, and temperature. conditions, followed by a 10 ns production run. Details are provided in Appendix F.
Stability.
Dynamic stability were analyzed as follows:
[…]
3.7 Protein-ligand Docking
Figure 4: Comparison of experimentally determined bindings (bold boxes) with predictions from blind docking simulations, using models closest to experimental data. Ligands are shown in green (experimental data), cyan (predictions from experimental receptors), magenta (predictions from SM samples), and yellow (predictions from FM samples). $\Delta G$ is in kcal/mol, with RMSD indicating deviation from experimental ligand positions (in Å). Dockings labeled “MD” represent protein-ligand complex MD simulations after blind docking.
Using AutoDock Vina (Trott & Olson, 2009; Eberhardt et al., 2021), we predicted optimal binding modes between generated structures and their family-specific ligands. We performed blind docking, scanning the entire protein surface for potential ligand binding sites without prior knowledge of binding pockets. The grid box covered the entire protein, and identical configurations were applied to both generated and experimental structures to evaluate if deviations in generated samples fell within an acceptable range, thus assessing their functional viability. Details on simulation settings are provided in Appendix G.
[…]
$\beta$-lactamases binding to penicillin.
Mutations at Glu166 and Asn170 in class A $\beta$-lactamases (Figure 1C) can form a stable acyl-enzyme intermediate, disrupting deacylation (Chen & Herzberg, 2001). Only asparagine and glycine at position 170 preserve WT-like function (Brown et al., 2009), and this conservation is retained in generated structures like FM-4 (Figure 4A), suggesting they may retain the ability to inactivate penicillin antibiotics.
Cytrochrome c binding to heme c.
The generated cytochrome c-like structures retain conserved residues found in the WT (Figure 1D; 1HRC). SM-1 (Figure 4B) includes phosphorylatable residues Tyr58, Thr59, and Tyr107, as well as lysine residues Lys82, Lys83, and Lys96 involved in phospholipid binding. Asn80 and these lysines form an ATP-binding pocket-like structure. FM-0 shows heme iron coordinated by two cysteines, which may form stronger covalent bonds, potentially affecting electron transfer efficiency.
KRas binding to GNP/GDP.
GNP, a non-hydrolyzable GTP analog, is commonly used to simulate the GTP-bound active state of Ras proteins (Pantsar, 2020). In WT KRas bound to GNP (Figure 4D; 5UFE), residues from the Switch II region fill the pocket. Removal of the $\gamma$-phosphate relaxes the protein conformation to the GDP-bound state, opening the switch II region (Figure 4C; 4OBE) (Kauke et al., 2017).
After blind docking the generated KRas-like structures with GNP/GDP, we analyzed the complexes using protein–ligand MD simulations. Allowing flexibility in both the backbone and side-chains enabled us to capture the dynamic conformational changes, especially in the switch regions, between active/inactive states. FM-generated KRas-like structures show greater flexibility. In FM-13 (Figure 4D), GNP binding causes switch II residues to fill the pocket, similar to 5UFE. In the GDP-bound state, FM-13 adopts an open conformation, with the channel between the switches and P-loop opening, as seen in 4OBE (Figure 4C). In contrast, SM-27 are more rigid, with fewer conformational changes between GDP- and GTP-bound states.
4 Discussion
This study demonstrates the potential of SM and FM methods for generating protein backbones. After determining the sequence, we performed side-chain homology modeling. Despite sequence dissimilarities, key residues are conserved, and the generated structures are highly similar. SM better captures conserved regions, producing more rigid structures, while FM offers greater flexibility. Structural phylogenetic analysis revealed evolutionary relationships, and MD and docking simulations confirmed stability and functionality. The presented pipeline and evaluation protocols build confidence for broader applications in protein generation.
SM and FM can generate a range of monomeric protein structures and can have applications beyond protein design (Yim et al., 2023; Bose et al., 2023). However, the complexity of these tasks is often underestimated, and function verification for novel samples remains costly. Targeting specific tasks or integrating generative methods into well-established empirical knowledge may yield better results. Key factors like conformational dynamics, water interactions, and entropy have not been fully considered in generation (Du et al., 2024). Incorporating protein sequences, side-chain details, and functional annotations as context could improve model performance. Although some progress has been made (Torge et al., 2023; Jin et al., 2023; Somnath et al., 2023; Zhou et al., 2023), research gaps remain.
These generative processes differ from protein synthesis simulations and do not reflect human hierarchical understanding. Limited interpretability reduces their utility for further downstream analyses and industrial applications. Although generated proteins may resemble real ones and include similar conserved residues and conformational changes, small variations could significantly impact function. Simply aiming for diversity and novelty does not guarantee quality.
Generative models have also shown great potential as efficient surrogates for MD simulations, which are often computationally intensive for complex systems or long-term behaviors. Jing et al. (2024) developed a framework to simulate molecular trajectories using generative models. Similarly, Viguera Diez et al. (2023) used generative models to enhance sampling of slow degrees of freedom, covering the sample space more effectively than traditional MD.
Software and Data
All input data and software used in this study are available from public sources or provided under academic licenses. The source code, scripts, generated samples, and curated datasets can be accessed at https://github.com/ECburx/PROTEVAL. The PDB structure files were downloaded on June 19, 2024, from https://www.wwpdb.org/ftp/pdb-ftp-sites.
Acknowledgments
The authors would like to thank Charlie Harris and Gos Micklem for their fruitful discussions and valuable feedback on this work.
Appendix A Protein Backbone Generation
A.1 $\mathrm{SE}(3)$ Decomposition into $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$
The Special Euclidean group $\mathrm{SE}(3)$ describes the rotations and translations in 3D space. An element of the $\mathrm{SE}(3)$ can be represented by a $4\times 4$ matrix:
$\bm{\mathsf{T}}=\begin{pmatrix}\bm{\mathsf{R}}&\bm{\mathsf{x}}\\ \mathbf{0}_{1\times 3}&1\end{pmatrix}$ (2)
where $\bm{\mathsf{R}}$ is a $3\times 3$ rotation matrix belonging to the Special Orthogonal group $\mathrm{SO}(3)$, and $\bm{\mathsf{x}}=[\mathsf{x}_{x}\ \mathsf{x}_{y}\ \mathsf{x}_{z}]\in\mathbb{R}^{3}$ is the translational component. Since $\mathrm{SE}(3)$ can be viewed as the semidirect product of $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$, denoted as $\mathrm{SE}(3)\cong\mathrm{SO}(3)\ltimes\mathbb{R}^{3}$, one option is to naturally treat $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$ as independent for simplicity (Yim et al., 2023).
A.2 Protein Backbone Representations
Molecules can be intuitively represented as 3D atomic point clouds. However, macromolecules like proteins may contain thousands or tens of thousands of atoms, with variation in the atom types and quantities among different amino acids (for instance, sulfur atoms are present only in a few amino acids like cysteine). Representing proteins as unordered 3D atomic point clouds significantly increases data dimensionality and sparsity, requiring far more training data than is typically available.
Following the work of Yim et al. (2023) and Bose et al. (2023), we adopt the more compact backbone rigid groups from AlphaFold (Jumper et al., 2021) to represent protein backbone structures in 3D space. A backbone rigid group consists of the main chain atoms (N, Cα, C, O) within a single residue (Figure 1A), where the their geometric relationships (relative positions and orientations) are highly stable. The position and orientation of the group is transformed as a whole, without accounting for individual atomic movements, simplifying the computation and reducing structural errors caused by excessive degrees of freedom.
Assuming experimentally measured ideal chemical bond angles and lengths, models learn how the rigid transformation (or frame) $\bm{\mathsf{T}}_{i}$ of each residue $i\in[1,N]$ acts on idealized coordinates $[\mathrm{N}^{\star},\mathrm{C}_{\alpha}^{\star},\mathrm{C}^{\star}]\in\mathbb{R}^{3}$ (centered at $\mathrm{C}_{\alpha}^{\star}$), so that the transformed coordinates match the actual coordinates as closely as possible:
$[\mathrm{N},\mathrm{C}_{\alpha},\mathrm{C}]_{i}=\bm{\mathsf{T}}_{i}\cdot[\mathrm{N}^{\star},\mathrm{C}_{\alpha}^{\star},\mathrm{C}^{\star}]$ (3)
where $\bm{\mathsf{T}}_{i}\in\mathrm{SE}(3)$ can be decomposed into a rotation matrix $\bm{\mathsf{R}}_{i}\in\mathrm{SO}(3)$ and a translation vector $\bm{\mathsf{x}}_{i}\in\mathbb{R}^{3}$. An additional torsion angle $\psi_{i}\in\mathrm{SO}(2)$ is introduced between the bond of $\mathrm{C}_{\alpha}$ and C for a more accurate construction of the backbone oxygen atom O.
A.3 $\mathrm{SE}(3)$ Score Matching
Let $\mathbf{T}_{t}=\left[\bm{\mathsf{T}}_{1,t},\dots,\bm{\mathsf{T}}_{N,t}\right]\in\mathrm{SE}(3)^{N}$ denote the manifold of $N$ frames at time $t$, where each frame can independently rotate and translate. Correspondingly, define $\mathbf{R}_{t}=\left[\bm{\mathsf{R}}_{1,t},\dots,\bm{\mathsf{R}}_{N,t}\right]$ and $\mathbf{X}_{t}=\left[\bm{\mathsf{x}}_{1,t},\dots,\bm{\mathsf{x}}_{N,t}\right]$. By treating $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$ as two independent stochastic processes, a forward process gradually perturbs the initial data distribution $p_{0}$. Following the approach of Yim et al. (2023), this process is described by the Stochastic Differential Equation (SDE) for $\mathbf{T}_{t}\sim p_{t}$ and any arbitrary time $t\in[0,T]$:
$\mathrm{d}\mathbf{T}_{t}=\left[0,-\frac{1}{2}\mathbf{X}_{t}\right]\mathrm{d}t+\left[\mathrm{d}\mathbf{B}_{t}^{\mathrm{SO}(3)},\mathrm{d}\mathbf{B}_{t}^{\mathbb{R}^{3}}\right]$ (4)
where $\mathbf{B}_{t}^{\mathrm{SO}(3)}$ and $\mathbf{B}_{t}^{\mathbb{R}^{3}}$ are Brownian motions on the $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$, respectively. Invariant density $p_{\mathrm{inv}}^{\mathrm{SE}(3)}(\bm{\mathsf{T}})\propto\mathcal{U}^{\mathrm{SO}(3)}(\bm{\mathsf{R}})\ \mathcal{N}(\bm{\mathsf{x}};0,\mathbf{I})$ is chosen for $\bm{\mathsf{T}}=(\bm{\mathsf{R}},\bm{\mathsf{x}})$.
Let $(\overleftarrow{\mathbf{T}}_{t})_{t\in[0,T]}=(\mathbf{T}_{T-t})_{t\in[0,T]}$, the corresponding time-reverse process (De Bortoli et al., 2022) is given by
$\mathrm{d}\overleftarrow{\mathbf{R}}_{t}$ $=\nabla_{\bm{\mathsf{R}}}\log p_{T-t}(\overleftarrow{\mathbf{T}}_{t})\mathrm{d}t+\mathrm{d}\mathbf{B}_{t}^{\mathrm{SO}(3)}$ (5)
$\mathrm{d}\overleftarrow{\mathbf{X}}_{t}$ =$=\left\{\frac{\overleftarrow{\mathbf{X}}_{t}}{2}+\nabla_{\bm{\mathsf{x}}}\log p_{T-t}(\overleftarrow{\mathbf{T}}_{t})\right\}\mathrm{d}t+\mathrm{d}\mathbf{B}_{t}^{\mathbb{R}^{3}}$ (6)
where $\nabla\log p$ is the gradient of the log-probability density function (also known as the Stein score). However, this gradient is typically intractable in practice because the exact form of $p_{t}(\mathbf{T}_{t})$ is unknown at any given time $t$.
Instead, score-based models estimates tractable conditional score $\nabla\log p_{t|0}$ through SM (Vincent, 2011), using a neural network $s(\theta,t,\cdot)$ trained by minimizing both [^4] One adds weights $1/\mathbb{E}[\lVert\nabla_{\mathbf{R}}\log p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})\rVert^{2}]$ to Equation 7 and $(1-e^{-t})/e^{-t/2}$ to Equation 8 for simplicity.:
$\mathcal{L}_{\text{SM}}^{\mathbf{R}}(\theta)$ $=\mathbb{E}\left[\lVert\nabla_{\mathbf{R}}\log p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})-s(\theta,t,\mathbf{R}_{t})\rVert^{2}\right]$ (7)
$\mathcal{L}_{\text{SM}}^{\mathbf{X}}(\theta)$ $=\mathbb{E}\left[\lVert\nabla_{\mathbf{X}}\log p_{t|0}(\mathbf{X}_{t}|\mathbf{X}_{0})-s(\theta,t,\mathbf{X}_{t})\rVert^{2}\right]$ (8)
with $t\sim\mathcal{U}(0,T)$ and
$\nabla_{\bm{\mathsf{R}}}\log p_{t|0}(\bm{\mathsf{R}}_{t}|\bm{\mathsf{R}}_{0})$ $=\frac{\bm{\mathsf{R}}_{t}}{\omega(\bm{\mathsf{R}}_{0\to t})}\log\{\bm{\mathsf{R}}_{0\to t}\}\frac{\partial_{\omega}f(\omega(\bm{\mathsf{R}}_{0\to t}),t)}{f(\omega(\bm{\mathsf{R}}_{0\to t}),t)}$ (9)
$\nabla_{\bm{\mathsf{x}}}\log p_{t|0}(\bm{\mathsf{x}}_{t}|\bm{\mathsf{x}}_{0})$ $=\frac{e^{-t/2}\bm{\mathsf{x}}_{0}-\bm{\mathsf{x}}_{t}}{1-e^{-t}}$ (10)
where $\omega$ represents the rotation angle, $\bm{\mathsf{R}}_{0\to t}=\bm{\mathsf{R}}_{0}^{\top}\bm{\mathsf{R}}_{t}$, and
$f(\omega,t)=\sum_{\ell\in\mathbb{N}}(2\ell+1)e^{-\ell(\ell+1)t/2}\frac{\sin((\ell+\frac{1}{2})\omega)}{\sin(\frac{\omega}{2})}$ (11)
is an auxiliary function for the heat kernel [^5] The heat kernel on a manifold is the fundamental solution to the heat equation, representing the probability density function of a Brownian particle diffusing from one point to another over time. of the Brownian motion on $\mathrm{SO}(3)$.
A.4 $\mathrm{SE}(3)$ Flow Matching
FM is a simulation-free method for training vector fields to follow a prescribed conditional probability path (Lipman et al., 2022). Formally, for $t\in[0,1]$, let $\mathbf{U}=\{\mathbf{u}_{t}\}$ be a flow which is a set of time-indexed vector fields that describe the paths along which data points move from an initial distribution $p_{1}$ to a target distribution $p_{0}$. Each vector field $\mathbf{u}_{t}(\mathbf{T}_{t})$ represents the rate of change of $\mathbf{T}_{t}$ which is typically the solution to the Ordinary Differential Equation (ODE) $\frac{d}{dt}\mathbf{T}_{t}=\mathbf{u}_{t}(\mathbf{T}_{t})$. FM approximates $\mathbf{u}_{t}(\mathbf{T}_{t})$ with a network $v(\theta,t,\cdot)$ by minimizing $\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}\left\lVert\mathbf{u}_{t}(\mathbf{T}_{t})-v(\theta,t,\mathbf{T}_{t})\right\rVert^{2}$ with $t\sim\mathcal{U}(0,1)$.
Similarly, independent flows can be built on $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$. Computing $\mathbf{u}_{t}$, however, is also intractable due to the complex integrals involved in defining the marginal probability path and vector field. By showing $\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta)$, Lipman et al. (2022) suggested the tractable conditional FM objective on $\mathbb{R}^{3}$:
$\mathcal{L}_{\text{CFM}}^{\mathbf{X}}(\theta)=\mathbb{E}\left\lVert\mathbf{u}_{t}(\mathbf{X}_{t}|\mathbf{X}_{0})-v(\theta,t,\mathbf{X}_{t})\right\rVert^{2}$ (12)
with the Gaussian path $p_{t}(\bm{\mathsf{x}}_{t}|\bm{\mathsf{x}}_{0})=\mathcal{N}(\bm{\mathsf{x}}_{t};t\bm{\mathsf{x}}_{0},(t\sigma-t+1)^{2})$ generated by
$\mathbf{u}_{t}(\bm{\mathsf{x}}_{t}|\bm{\mathsf{x}}_{0})=\frac{\bm{\mathsf{x}}_{0}-(1-\sigma)\bm{\mathsf{x}}_{t}}{1-(1-\sigma)t}$ (13)
where $\sigma>0$ is a smoothing constant.
For flows on $\mathrm{SO}(3)$, Bose et al. (2023) set
$\mathcal{L}_{\text{CFM}}^{\mathbf{R}}(\theta)=\mathbb{E}\left\lVert\mathbf{u}_{t}(\mathbf{R}_{t}|\mathbf{R}_{0},\mathbf{R}_{1})-\mathbf{v}(\theta,t,\mathbf{R}_{t})\right\rVert^{2}$ (14)
and define the geodesic interpolant between $\bm{\mathsf{R}}_{1}\sim p_{1}$ and $\bm{\mathsf{R}}_{0}\sim p_{0}$ as $\bm{\mathsf{R}}_{t}=\exp_{\bm{\mathsf{R}}_{1}}(t\log_{\bm{\mathsf{R}}_{1}}(\bm{\mathsf{R}}_{0}))$. Let $\Psi_{t}$ be a flow that connects $\bm{\mathsf{R}}_{1}$ to $\bm{\mathsf{R}}_{0}$, computing $\mathbf{u}_{t}(\bm{\mathsf{R}}_{t}|\bm{\mathsf{R}}_{0},\bm{\mathsf{R}}_{1})$ simplifies to determining $\bm{\mathsf{R}}_{t}$ along $\frac{\mathrm{d}}{\mathrm{d}t}\Psi_{t}(\bm{\mathsf{R}})=\dot{\bm{\mathsf{R}}}_{t}$ (Chen & Lipman, 2023) and then taking its time derivative. Thus, we have
$\mathbf{u}_{t}(\bm{\mathsf{R}}_{t}|\bm{\mathsf{R}}_{0},\bm{\mathsf{R}}_{1})=\frac{\log_{\bm{\mathsf{R}}_{t}}(\bm{\mathsf{R}}_{0})}{t}$ (15)
Optimal transport.
Optimal Transport (OT) conditions hold when the probability paths between two distributions are defined by a displacement map that linearly interpolates between them (Pooladian et al., 2023).
Tong et al. (2023) views the OT problem as finding a mapping that minimizes the 2-Wasserstein distance between two distributions $p_{1}$ and $p_{0}$ on $\mathbb{R}^{3}$, using the Euclidean distance $\lVert\bm{\mathsf{x}}_{0}-\bm{\mathsf{x}}_{1}\rVert$ as the displacement cost:
$W(p_{0},p_{1})^{2}_{2}=\inf_{\pi\in\Pi}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\lVert\bm{\mathsf{x}}_{0}-\bm{\mathsf{x}}_{1}\rVert^{2}\mathrm{d}\pi(\bm{\mathsf{x}}_{0},\bm{\mathsf{x}}_{1})$ (16)
where $\Pi$ denotes the set of all joint probability measures on $\mathbb{R}^{3}\times\mathbb{R}^{3}$ with marginals $p_{1}$ and $p_{0}$. By setting $p(\bm{\mathsf{x}}_{0},\bm{\mathsf{x}}_{1})=\pi(\bm{\mathsf{x}}_{0},\bm{\mathsf{x}}_{1})$ and a Gaussian conditional probability path with mean $\mu_{t}=t\bm{\mathsf{x}}_{0}+(1-t)\bm{\mathsf{x}}_{1}$, we have
$\mathcal{L}_{\text{OT}}^{\mathbf{X}}(\theta)=\mathbb{E}_{\pi}\left\lVert\mathbf{u}_{t}(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{X}_{1})-v(\theta,t,\mathbf{X}_{t})\right\rVert^{2}$ (17)
$\mathbf{u}_{t}(\bm{\mathsf{x}}_{t}|\bm{\mathsf{x}}_{0},\bm{\mathsf{x}}_{1})=\bm{\mathsf{x}}_{0}-\bm{\mathsf{x}}_{1}$ (18)
with $p_{t}(\bm{\mathsf{x}}_{t})=\int\mathcal{N}(\bm{\mathsf{x}}_{t}|t\bm{\mathsf{x}}_{0}+(1-t)\bm{\mathsf{x}}_{1},\sigma^{2})\pi(\bm{\mathsf{x}}_{0},\bm{\mathsf{x}}_{1})\mathrm{d}\bm{\mathsf{x}}_{0}\mathrm{d}\bm{\mathsf{x}}_{1}$.
Inspired by this, Bose et al. (2023) extended Equation 14 and Equation 15 to $\mathrm{SO}(3)$ using Riemannian optimal transport, with $\bar{\pi}$ being the projection of $\pi$ on $\mathrm{SO}(3)$:
$\mathcal{L}_{\text{OT}}^{\mathbf{R}}(\theta)=\mathbb{E}_{\bar{\pi}}\left\lVert\frac{\log_{\mathbf{R}_{t}}(\mathbf{R}_{0})}{t}-v(\theta,t,\mathbf{R}_{t})\right\rVert^{2}$ (19)
A.5 $\mathrm{SE}(3)$ Invariance
$\mathrm{SE}(3)$ invariance can be achieved by consistently positioning the model at the origin (Yim et al., 2023; Bose et al., 2023; Rudolph et al., 2020).
In the context of SM, to ensure translation invariance on $\mathbb{R}^{3}$, one apply a projection matrix $\mathbf{P}\in\mathbb{R}^{3N\times 3N}$ that removes the center of mass $\frac{1}{N}\sum_{i=1}^{N}\bm{\mathsf{x}}_{i}$. It results in an invariant measure on $\mathrm{SE}(3)^{N}$, denoted as $\mathrm{SE}(3)^{N}_{0}$. Since the Brownian motion on $\mathrm{SO}(3)$ and the score $\nabla_{\bm{\mathsf{R}}}\log p_{T-t}$ are both rotation-invariant, Equation 5 is $\mathrm{SO}(3)$-invariant. Consequently, Yim et al. (2023) derive the following $\mathrm{SE}(3)$-invariant forward process
$\mathrm{d}\mathbf{T}_{t}=\left[0,-\frac{1}{2}\mathbf{P}\mathbf{X}_{t}\right]\mathrm{d}t+\left[\mathrm{d}\mathbf{B}_{t}^{\mathrm{SO}(3)^{N}},\mathbf{P}\mathrm{d}\mathbf{B}_{t}^{\mathbb{R}^{3N}}\right]$ (20)
and its corresponding time-reverse process
$\mathrm{d}\overleftarrow{\mathbf{R}}_{t}$ $=\nabla_{\bm{\mathsf{R}}}\log p_{T-t}(\overleftarrow{\mathbf{T}}_{t})\mathrm{d}t+\mathrm{d}\mathbf{B}_{t}^{\mathrm{SO}(3)^{N}}$ (21)
$\mathrm{d}\overleftarrow{\mathbf{X}}_{t}$ $=\mathbf{P}\left\{\frac{\overleftarrow{\mathbf{X}}_{t}}{2}+\nabla_{\bm{\mathsf{x}}}\log p_{T-t}(\overleftarrow{\mathbf{T}}_{t})\right\}\mathrm{d}t+\mathrm{d}\mathbf{P}\mathbf{B}_{t}^{\mathbb{R}^{3N}}$ (22)
The same approach can be applied to FM. After centering and decoupling the flow on $\mathrm{SE}(3)_{0}^{N}$, a separate $\mathrm{SE}(3)$-invariant flow can be constructed for each residue in backbone [^6] As the product group of $N$ copies of $\mathrm{SE}(3)$, $\mathrm{SE}(3)_{0}^{N}$ has a geometric structure that allows global geometric operations (such as geodesic distance, exponential maps, and logarithmic maps) to be decomposed into operations on each of the $N$ $\mathrm{SE}(3)$ groups., in which each $\mathrm{SE}(3)$-invariant measure is decomposed into a measure that is proportional to the Lebesgue measure on $\mathbb{R}^{3}$ (Pollard, 2001) and an $\mathrm{SO}(3)$-invariant measure (Bose et al., 2023).
A.6 Additional Losses
To prevent unrealistic fine-grained features such as steric clashes or chain breaks when learning the torsion angle $\psi$, Yim et al. (2023) proposed adding two additional loss functions. The first is the mean squared error (MSE) on backbone atom positions:
$\mathcal{L}_{\text{bb}}=\frac{1}{4N}\sum_{n=1}^{N}\sum_{a\in A}\lVert a_{n}-\hat{a}_{n}\rVert^{2}$ (23)
where $A=\{\mathrm{N},\,\mathrm{C},\,\mathrm{C}\alpha,\,\mathrm{O}\}$. $a_{n}$ and $\hat{a}_{n}$ are the true and predicted coordinates of backbone atom $a$ at residue $n$.
The second loss penalizes deviations in local pairwise atomic distances:
[…]
where $d_{ab}^{nm}=\lVert a_{n}-b_{m}\rVert$ and $\hat{d}_{ab}^{nm}$ are the true and predicted distances between atoms $a$ and $b$ in residues $n$ and $m$, respectively. The indicator function $\mathds{1}\{d_{ab}^{nm}<6\text{\AA}\}$ limits the loss to atom pairs within 6 Å̊$\mathring{\mathrm{A}}$.
The complete training loss is given by
$\mathcal{L}=\mathcal{L}^{\mathbf{R}}(\theta)+\mathcal{L}^{\mathbf{X}}(\theta)+\mathds{1}\left\{t<T/4\right\}\left(\mathcal{L}_{\text{bb}}+\mathcal{L}_{\text{2D}}\right)$ (25)
where $T=1$ in the case of FM.
A.7 Model Architecture
Figure 5: Overview of the (A) embedding module and (B) multi-layer network architecture.
The networks $s(\theta,t,\cdot)$ involved in SM and $v(\theta,t,\cdot)$ involved in FM models, as reviewed in Section 2, can share a common high-level architecture (Yim et al., 2023; Bose et al., 2023; Anand & Achim, 2022).
Embeddings.
Given node embedding dimensions $D_{h}$ and edge embedding dimensions $D_{e}$, node embeddings $\mathbf{h}\in\mathbb{R}^{N\times D_{h}}$ are derived from residue indices $\textbf{i}=\{1,\dots,N\}$ and time-step information $\textbf{t}=\{0,\Delta t,\dots,T\}$, while edge embeddings $\mathbf{E}=\in\mathbb{R}^{N\times N\times D_{e}}$ integrate additional features, such as relative sequence distances $j-i$ for any $i,j\in[1,N]$ (Figure 5A). Self-conditioning on the predicted $\mathrm{C}_{\alpha}$ displacements is also applied:
$c_{i,j}=\sum_{b=1}^{B}\mathds{1}\left\{|\bm{\mathsf{x}}_{i}^{*}-\bm{\mathsf{x}}_{j}^{*}|<v_{b}\right\}$ (26)
where $\bm{\mathsf{x}}^{*}$ denotes the coordinates for $\mathrm{C}_{\alpha}$ predicted through self-conditioning, and $v_{1},\dots,v_{B}$ are bins spaced uniformly from 0 and $B$ angstroms. These initial features are encoded using multilayer perceptrons (MLPs) along with sinusoidal embeddings (Vaswani et al., 2017).
Multi-layer network.
Figure 5B shows the architecture of the multi-layer neural network ($L=4$ layers used in our experiments). At each layer $l$, the network takes node embeddings $\textbf{h}^{l}$, edge embeddings $\textbf{E}^{l}$, and rigid transformations $\textbf{T}^{l}$ as input, applying the Invariant Point Attention (IPA) introduced by Jumper et al. (2021) to enable spatial attention. Transformer from Vaswani et al. (2017) models interactions along the chain structure. The network’s update procedure remains invariant under $\mathrm{SE}(3)$ transformations due to the inherent $\mathrm{SE}(3)$-invariance of the IPA.
The output $\textbf{T}^{L}$ from the final layer serves as the predicted frame, denoted as $\hat{\textbf{T}}_{0}=(\hat{\textbf{R}}_{0},\hat{\textbf{x}}_{0})$. Consequently, for SM, we have the following scores predictions based on Equation 9 and Equation 10:
$\forall s(\theta,t,\bm{\mathsf{R}}_{t})\in s(\theta,t,\mathbf{R}_{t}),\quad s(\theta,t,\bm{\mathsf{R}}_{t})$ $=\nabla_{\bm{\mathsf{R}}}\log p_{t|0}(\bm{\mathsf{R}}_{t}|\hat{\bm{\mathsf{R}}}_{0})$ (27)
$=\frac{\bm{\mathsf{R}}_{t}}{\omega(\hat{\bm{\mathsf{R}}}_{0}^{\top}\bm{\mathsf{R}}_{t})}\log\{\hat{\bm{\mathsf{R}}}_{0}^{\top}\bm{\mathsf{R}}_{t}\}\frac{\partial_{\omega}f(\omega(\bm{\mathsf{R}}_{0}^{\top}\bm{\mathsf{R}}_{t}),t)}{f(\omega(\bm{\mathsf{R}}_{0}^{\top}\bm{\mathsf{R}}_{t}),t)}$ (28)
$\forall s(\theta,t,\bm{\mathsf{x}}_{t})\in s(\theta,t,\mathbf{X}_{t}),\quad s(\theta,t,\bm{\mathsf{x}}_{t})$ $=\nabla_{\bm{\mathsf{x}}}\log p_{t|0}(\bm{\mathsf{x}}_{t}|\hat{\bm{\mathsf{x}}}_{0})$ (29)
$=\frac{e^{-t/2}\hat{\bm{\mathsf{x}}}_{0}-\bm{\mathsf{x}}_{t}}{1-e^{-t}}$ (30)
From Equation 13 and Equation 15, we have the following for FM with OT:
$\forall v(\theta,t,\bm{\mathsf{R}}_{t})\in v(\theta,t,\mathbf{R}_{t}),\quad v(\theta,t,\bm{\mathsf{R}}_{t})$ $=\mathbf{u}_{t}(\bm{\mathsf{R}}_{t}|\hat{\bm{\mathsf{R}}}_{0},\bm{\mathsf{R}}_{1})$ (31)
$=\frac{\log_{\bm{\mathsf{R}}_{t}}(\hat{\bm{\mathsf{R}}}_{0})}{t}$ (32)
$\forall v(\theta,t,\bm{\mathsf{x}}_{t})\in v(\theta,t,\mathbf{X}_{t}),\quad v(\theta,t,\bm{\mathsf{x}}_{t})$ $=\mathbf{u}_{t}(\bm{\mathsf{x}}_{t}|\hat{\bm{\mathsf{x}}}_{0})$ (33)
$=\frac{\hat{\bm{\mathsf{x}}}_{0}-(1-\sigma)\bm{\mathsf{x}}_{t}}{1-(1-\sigma)t}$ (34)
Torsion angle $\bm{\hat{\psi}}=\{\hat{\psi}_{1},\dots,\hat{\psi}_{N}\}=\bm{\psi}/\lVert\bm{\psi}\rVert\in\mathrm{SO(2)}^{N}$ is predicted with $\textbf{h}^{L}$ and $\textbf{E}^{L}$.
Appendix B Amino Acid Count Distributions in Experimental Proteins Used for Training
Figure 6: (A-D) Distribution of amino acid sequence lengths (aa) in experimentally determined protein structures used for training.
Appendix C Structural Phylogenetic Trees
Figure 7 shows structural phylogenetic trees generated using $Q_{\text{score}}$ and the 3Di alphabet. Both approaches reveal evolutionary relationships between the generated and experimental structures. However, compared to the $Q_{\text{score}}$-based tree, the 3Di alphabet tree shows closer and less distinguishable evolutionary relationships among samples from different protein families, especially between the GFP and cytochrome c samples.
Figure 7: Structural phylogenetic trees constructed using (A) $Q_{\text{score}}$ and (B) 3Di alphabet. Members of different protein families are grouped into semi-circular rings. The colored branches and nodes represent the generated protein structures.
Appendix D Conserved Residue Consistency
Figure 8: Normalized evolutionary rates were mapped onto sequences, with black indicating rapidly evolving positions and orange indicating conserved positions. The heat map displays each column as the optimal predicted sequence of a generated sample, with rows representing specific positions in these sequences. The bar chart (exp. identity) shows the average pairwise sequence identity at each position from the MSAs of experimental proteins. Taller bars signify positions where most experimental protein share the same residue. Consistency between generated and experimental sequences at conserved positions appears as uniform orange coloring in the heat map, aligning with the taller bars.
Appendix E Evaluation of Side-Chain Homology Modeling
We used homology modeling to add side-chains to the generated protein backbones, evaluating them using PROCHECK (Laskowski et al., 1993, 1996) and WHAT_CHECK (Hooft et al., 1996) to correct or exclude those not meeting the expectations. Specifically:
Planarity.
Planar side-chains, such as those in phenylalanine, tyrosine, tryptophan, and histidine, are essential for stability and function. Conformations lacking expected planarity were discarded.
Asparagine, glutamine, histidine flips.
Asparagine, glutamine, and histidine side-chains can experience terminal flips, altering key interactions. WHAT_CHECK was used to evaluate and, if necessary, adjust side-chain orientations to have more stable interactions.
Torsion angles.
Side-chain torsion angles ($\chi$ angles) were assessed, focusing on $\chi_{1}$ (rotation around $\mathrm{C}_{\alpha}$ to the first side-chain atom) and $\chi_{2}$ (rotation to the second side-chain atom) to prevent spatial clashes. Conformations in uncommon $\chi$-angle regions were excluded.
Bond lengths and angles.
Unusual bond lengths and angles may indicate strain and modeling errors, potentially disrupting interactions. Conformations with such issues were discarded.
Other parameters.
Side-chains with abnormal torsion angles, atypical aromatic bonding angles, or unusual proline puckering were also discarded.
Appendix F Molecular Dynamics Simulations
For experimental structures, crystallographic water and unnecessary small molecules were removed. For generated structures, missing side-chains were added via homology modeling (Section 3.6). After adding hydrogen atoms using Reduce2 (Grosse-Kunstleve et al., 2002) and confirming no missing atoms, each protein was centered at the origin.
Simulations were performed with GROMACS (Abraham et al., 2015), using the all-atom CHARMM36 force field (July 2022 version) (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Vanommeslaeghe et al., 2012; Yu et al., 2012; Soteras Gutiérrez et al., 2016).
F.1 Molecular Dynamics Setup for Stability Assessment of Generated Structures
[…]
In accordance with GROMACS 2024 documentation, we applied the following configurations in the MD parameter (.mdp) files. Van der Waals interactions were handled using a cutoff method, while long-range electrostatic interactions were calculated using the Particle Mesh Ewald (PME) method.
constraints     = h-bonds
cutoff-scheme   = Verlet
vdwtype         = cutoff
vdw-modifier    = force-switch
rlist           = 1.2
rvdw            = 1.2
rvdw-switch     = 1.0
coulombtype     = PME
rcoulomb        = 1.2
DispCorr        = no
After solvating the system with water using the TIP3P model, we added Na+ and Cl- ions to achieve a physiological concentration of 150 mM and to neutralize the system’s total charge. Energy minimization was then conducted to resolve steric clashes and optimize the geometry, with potential energy and maximum force monitored to ensure they reached acceptable thresholds.
The next step involved equilibrating the solvent and ions around the protein. We chose the leap-frog integrator for the simulations and applied the LINCS algorithm to constrain hydrogen bonds. Equilibration involved two stages. In the first stage, we performed a 500 ps NVT equilibration (250,000 steps with a 2 fs time step). Temperature control was managed using the V-rescale thermostat, with the system divided into two groups: (1) protein and (2) water + ions, both set to a target temperature of 310 K to simulate physiological conditions. In the second stage, we carried out a 500 ps NPT equilibration with pressure coupling enabled. The pressure was regulated using the C-rescale method with isotropic coupling. The target pressure was 1.0 bar, with a compressibility of $4.5\times 10^{-5}$ bar-1 and a pressure coupling time constant of 0.5 ps.
Following equilibration, we conducted a 10 ns production simulation (5,000,000 steps with a 2 fs time step), during which all position restraints were removed. This allowed us to observe and analyze the system’s dynamic behavior over time, in order to access its stability. Full details of the MD parameter files can be found in Software and Data.
F.2 Molecular Dynamics Setup for Conformational Analysis of Protein-Ligand Complexes
The receptor and ligand were saved as separate coordinate files to prepare their respective topologies. The receptor topology was prepared as in Section F.1. For the ligand, hydrogen atoms were added using OpenBabel (O’Boyle et al., 2011), and topology was generated via the CGenFF server (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Vanommeslaeghe et al., 2012). The receptor and ligand topologies, along with force-field-compatible coordinate files, were then combined to construct the complete complex system.
The MD workflow for complexes followed the same steps in Section F.1. Complexes were placed in an octahedral simulation box, energy-minimized in vacuum, solvated in water, and neutralized with Na+ and Cl- ions to 150 mM. A second energy minimization was then performed on the solvated system.
During equilibration, positional restraints were applied to the ligand to prevent unnecessary displacement in the initial stages of the simulation. Additionally, to minimize interference from temperature fluctuations of the ligand on the overall simulation, we defined two temperature coupling groups: (1) the receptor and ligand as one group, and (2) the solvent and ions as the other. Other equilibration settings followed Section F.1.
After equilibration, restraints were removed, and a 10 ns production simulation was conducted to analyze the dynamic behavior and conformational changes in the complexes.
Appendix G Protein-ligand Blind Docking
Similarly, crystallographic water and unwanted molecules were removed from experimental structures, and missing side-chains were added to generated structures via homology modeling. Receptor structures were prepared using AutoDock Tools (Morris et al., 2009), with polar hydrogens added, Kollman charges assigned, and any missing atoms repaired. For receptors within the same family, we prepared a shared ligand file, adding hydrogen atoms and assigning Gasteiger charges. A large grid box, typically 80 to 110 Å̊$\mathring{\text{A}}$ per side, was defined to cover the entire protein surface.
Using these settings, we performed blind docking with AutoDock Vina (Trott & Olson, 2009; Eberhardt et al., 2021), generating up to 25 binding modes with a maximum energy difference of 5 kcal/mol and an exhaustiveness level of 20. The binding mode with the lowest binding free energy was selected as the final result.
Appendix H Generated Structures
Figure 9: 50 $\beta$-lactamase-like protein backbones generated using score matching and 50 using flow matching.
Figure 10: 50 cytochrome c-like protein backbones generated using score matching and 50 using flow matching.
Figure 11: 50 GDP-like protein backbones generated using score matching and 50 using flow matching.
Figure 12: 50 Ras-like protein backbones generated using score matching and 50 using flow matching.

### v2

@@
-Building Confidence in Deep Generative Protein Design
+Challenges and Guidelines in Deep Generative Protein Design: Four Case Studies
 Abstract
-Deep generative models show promise for de novo protein design, but their effectiveness within specific protein families remains underexplored.
-In this study, we evaluate two 3D rigid-body generative methods, score matching and flow matching, to generate monomeric protein backbones in $\mathrm{SE}(3)$ space.
-Our goal is to provide new insights and build confidence in the broader applicability of deep generative models for protein design.
-The optimal amino acid sequences were predicted from the generated backbones, followed by side-chain homology modeling.
-Results demonstrated high structural integrity, with conserved key residues aligning with known proteins.
-Structural phylogenetic analysis shows evolutionary links between the generated samples and their protein family members.
-Further molecular dynamics simulations and protein-ligand docking confirm the dynamic stability and functional potential of these samples, with ligand binding inducing conformational changes consistent with those in wild-type proteins.
-Machine Learning
+Deep generative models show promise for de novo protein design, yet reliably producing designs that are geometrically plausible, evolutionarily consistent, functionally relevant, and dynamically stable remains challenging.
+We present a deep generative modeling pipeline for early de novo design of monomeric proteins, based on Score Matching and Flow Matching.
+We apply this pipeline to four diverse protein families with an adaptable evaluation protocol.
+Generated structures display realistic, clash‐free conformations enriched with family-specific features, while the designed sequences preserve essential functional residues while retaining variability.
+Molecular dynamics and binding simulations show dynamic stability, with wild-type-like binding pockets that interact favorably with family-specific ligands.
+These results provide practical guidelines for integrating generative models into protein design workflows.
+Generative Protein Design, Score Matching, Flow Matching, Structural Phylogenetics, Molecular Dynamics, Ligand Docking
 1 Introduction
-Protein functions and specificity are determined by their complex structures.
-Over the past 60 years, we have progressed from viewing protein design as unattainable to achieving complete artificial design and synthesis (Korendovych & DeGrado, 2020; Huang et al., 2016), with expanding applications across industries (Arunachalam et al., 2021; Kingwell, 2024; Barclay & Acharya, 2023).
-However, precise design remains challenging due to the nonlinear complexity of protein folding and the sensitivity of function to slight changes, still demanding significant resources.
+Protein functions and specificities are dictated by their complex structures.
+Over the past 60 years, we have progressed from viewing protein design as unattainable to achieving complete artificial design and synthesis (Korendovych & DeGrado, 2020; Huang et al., 2016), with expanding applications across industries (Arunachalam et al., 2021; Kingwell, 2024; Barclay & Acharya, 2023; Victorino da Silva Amatto et al., 2021; Ali et al., 2020).
+However, achieving atomic-level precision remains challenging due to the nonlinear complexity of folding and the sensitivity of the function to slight changes, still demanding significant resources.
 With the rapid growth of protein structure databases (Jamasb et al., 2024; Berman, 2000), various in silico de novo protein design methods, particularly deep generative models (Watson et al., 2023; Ingraham et al., 2023; Wu et al., 2022), have emerged.
-Approaches such as diffusion-based score matching (SM) (Yim et al., 2023), and flow matching (FM) (Bose et al., 2023), show competitive performance for generating monomeric protein backbones in $\mathrm{SE}(3)$ [^1] Special Euclidean group.
-See more in Section A.1.
-space.
-While many methods show novelty, diversity, and designability in generated structures, few studies assess their functionality or evolutionary relevance, leaving the observed novelty and diversity potentially as mere hallucinations (Ji et al., 2023).
-Unclear biological functionality, unknown stability, and the lack of validation connecting these structures to known functions limit the broader application and advancement of generative methods.
-This study evaluates the effectiveness of SM and FM methods across four well-studied protein families ($\beta$-lactamases, cytochrome c, GFP, and Ras).
-After determining optimal amino acid sequences and conducting side-chain homology modeling, we assess the functional relevance and evolutionary potential of the design, aiming to build confidence in deep generative approaches for protein design.
-Our key contributions are: (1) We present a deep generative model pipeline for early-stage de novo protein design, with a robust evaluation protocol adaptable to a wider range of protein structure generation methods.
-(2) Despite sequence dissimilarity, generated samples show structural similarity to experimentally determined proteins in their families, conserving key residues linking to functions.
-(3) Structural phylogenetic analysis of both experimental and generated proteins reveals evolutionary relationships, showing the potential of structural phylogenetics.
-(4) Molecular dynamics (MD) simulations and protein-ligand docking show the dynamic stability and functional potential of generated samples.
-These samples form binding pockets akin to wild-type (WT) proteins, with conserved residues that bind family-specific ligands at low energies.
-For instance, generated KRas-like structures exhibit switch region flexibility and GTP/GDP-dependent conformational shifts similar to WT.
-(5) FM samples show greater flexibility, diversity, and novelty, while SM better captures conserved regions, offering insights for model selection and optimization in future work.
-2 Background: Protein Backbone Generation
-In this section and Appendix A, we review the key concepts behind Yim et al. (2023) and Bose et al. (2023)’s approaches for backbone generation, using SM and FM on $\mathrm{SE}(3)$.
-Protein backbone representations.
-Molecules can be intuitively represented as atomic point clouds; however, for macromolecules, this approach often leads to high dimensionality and data sparsity.
-A more compact, fragment-based representation involves using backbone rigid groups from AlphaFold (Jumper et al., 2021).
-For each amino acid residue $i\in[1,N]$, its group consists of the backbone atoms $[\mathrm{N},\mathrm{C}_{\alpha},\mathrm{C},\mathrm{O}]_{i}$ (Figure 1A).
-The goal is to learn how the rigid transformation (or frame) $\bm{\mathsf{T}}_{i}$ acts on the idealized coordinates, so that the transformed coordinates match the actual coordinates.
-For simplicity, $\bm{\mathsf{T}}_{i}\in\mathrm{SE}(3)$ decomposes into a rotation matrix $\bm{\mathsf{R}}_{i}\in\mathrm{SO}(3)$ and a translation vector $\bm{\mathsf{x}}_{i}\in\mathbb{R}^{3}$.
-An extra dihedral angle $\psi_{i}\in\mathrm{SO}(2)$ is introduced between the $\mathrm{C}_{\alpha}$ and C bond to optimize the positioning of the backbone oxygen ($\mathrm{O}$).
-$\mathrm{SE}(3)$ score matching.
-Let $\mathbf{T}_{t}=\left[\bm{\mathsf{T}}_{1,t},\dots,\bm{\mathsf{T}}_{N,t}\right]$ be the manifold of $N$ frames at time $t$.
-The forward process gradually perturbs the distribution $\mathbf{T}_{0}\sim p_{0}$, pushing it toward a simple distribution $p_{T}$, such as a Gaussian, by introducing a stochastic process independently in rotation and translation spaces (Equation 4).
-The reverse process approximately reconstructs $p_{0}$ by iteratively calculating the gradient of the log probability density (or the score) $\nabla\log p$ along a reverse path (Equation 5 and Equation 6).
-Since this score function is typically intractable in practice, a network $s(\theta,t,\cdot)$ is trained to approximate the conditional score $\nabla\log p_{t|0}$ at time $t\sim\mathcal{U}(0,T)$ (Vincent, 2011).
-$\mathrm{SE}(3)$ flow matching.
-The flow from $p_{T}$ to $p_{0}$ is defined as a set of continuously differentiable, time-indexed vector fields $\{\mathbf{u}_{t}\}$, where FM learns the time evolution of $\mathbf{u}_{t}$ to gradually transport the data from $p_{T}$ to $p_{0}$.
-Because $\mathbf{u}_{t}$ is generally intractable, Lipman et al.
-(2022) introduced conditional FM, training a neural network $v(\theta,t,\cdot)$ to approximate $\mathbf{u}_{t|0}$ (Equation 13).
-When the optimal transport condition is satisfied, it provides a smooth and efficient transformation path, improving training stability and reducing computational complexity (Pooladian et al., 2023).
-3 Evaluations of Generated Structures
-Figure 1: (A) Protein backbone and dihedral angles $\psi$ and $\phi$.
-(B) Ramachandran plots showing $\psi$ and $\phi$ angle distributions for generated protein backbones vs.
-experimental proteins.
-The top-right inset shows conformationally favored (light) and disallowed (dark) angle regions.
-(C) Structural overlay of WT E.
-coli TEM1 (PDB: 1BTL; Jelsch et al.
-(1993); green) and its E166N acylated intermediate (PDB: 1FQG; Brown et al.
-(2009); white) with penicillin (PNM).
-(D) Structure of the WT E.
-caballus heart cytochrome c (PDB: 1HRC; (Dickerson et al., 1967)) with heme C (HEC).
-(E) Structure of A.
-victoria GFP and its chromophore (PDB: 4KW4; Barnard et al.
-(2014)).
-(F) Structure of GDP-bound H.
-sapiens KRas protein (PDB: 4OBE; (Hunter et al., 2014)).
-3.1 Data
-We generated monomeric proteins across a range of families, covering diverse fold types and functions.
-These proteins include both natural and engineered mutations while preserving conserved core functional regions.
-$\beta$-lactamases.
-$\beta$-lactamases are enzymes that deactivate $\beta$-lactam antibiotics by hydrolyzing their $\beta$-lactam ring, contributing significantly to bacterial resistance (Lee et al., 2016).
-Using $\beta$-lactamase inhibitors can restore antibiotic efficacy by blocking this reaction (Behzadi et al., 2020).
-Since their discovery, $\beta$-lactamase diversity has rapidly expanded due to the evolution of bacterial resistance, making them ideal for protein modeling studies.
-For this study, we gathered structural data on 1,578 unique monomeric $\beta$-lactamases and their variants across Ambler classes (A, B, C, and D) from the BLDB (Naas et al., 2017) and the Protein Data Bank (PDB) (Berman, 2000).
-Class A $\beta$-lactamases are the most prevalent, with conserved active-site residues Ser70, Glu166, and Asn170 coordinating the hydrolytic water for deacylation (Tooke et al., 2019; Brown et al., 2009).
-Figure 1C shows the wild-type $\beta$-lactamase TEM1 (1BTL) alongside its acylated E166N intermediate (1FQG).
-Cytochrome c.
-Cytochrome c is a water-soluble protein ($\sim 12$ kDa) essential for ATP synthesis in mitochondria and intrinsic apoptosis (Kashyap et al., 2021; Ow et al., 2008).
-It also serves as an independent marker for apoptosis in several cancers (Li et al., 2001; Way et al., 2004).
-Despite variations across species, its core structure and function are conserved.
-We obtained structural data for 498 unique cytochrome c proteins and variants from the PDB.
-Figure 1D shows horse cytochrome c, where a hydrophobic shell surrounds the heme group, with only $\sim 7.5\%$ of the surface available for electron transfer (Bushnell et al., 1990).
-The hydrophobic environment and iron coordination by His18 and Met80 maintain a high redox potential ( 260 mV) (Salemme, 1977).
-Phosphorylation occurs at Thr28, Thr47, Tyr48, and Tyr97 (Hüttemann et al., 2011), while Lys72, Lys73, and Lys87 bind phospholipids (Kagan et al., 2009).
-The ATP-binding pocket includes Glu69, Asn70, Lys88, and Lys72, Lys86, Lys87 (McIntosh et al., 1996).
-Green fluorescent proteins (GFP).
-GFP, first isolated from Aequorea victoria, emits green fluorescence under specific wavelengths.
-Its core structure is an 11-strand $\beta$-barrel that encloses the chromophore (Figure 1E) (Remington, 2011).
-Various mutants have been engineered to enhance or modify its properties, including enhanced GFP (Cormack et al., 1996), superfolder GFP (Pédelacq et al., 2005), and color variants like YFP (Ormö et al., 1996) and BFP (Glaser et al., 2016).
-We collected structural data for 448 GFPs and variants from the PDB to explore this diversity.
-Ras.
-Ras proteins, a subgroup of the small GTPase superfamily, act as molecular switches, cycling between GTP-bound (active) and GDP-bound (inactive) states to regulate cell proliferation, differentiation, migration, and apoptosis (Ladygina et al., 2011; Simanshu et al., 2017; Weinmann & Ottow, 2007).
-They play a key role in signaling from the cell surface to downstream pathways.
-Mutations that keep Ras proteins in an active state drive excessive cell growth and malignancy, making Ras inhibition a promising cancer treatment strategy (Singh & Lingham, 2002).
-We focused on the most common cancer-related Ras proteins (HRas, KRas, and NRas) (Cox, 2002) and obtained 511 experimental structures from the PDB.
-In human KRas (Figure 1F), the switch I and II regions form the key interface for interacting with effectors and regulators (Pantsar, 2020).
-These regions are highly flexible, with conformations depending on GTP or GDP binding.
-Cancer-related mutations frequently occur in the P-loop and switch II (Pantsar, 2020).
-3.2 Training and Sampling
-We trained the SM and FM (with OT) models on the described protein families, using pretrained weights from Yim et al.
-(2023) and Bose et al.
-(2023), with each model having $\sim 17$ million parameters.
-The sequence length distributions of the training proteins are shown in Figure 6.
-Each model generated 50 backbone structures with target lengths of 270 amino acids for $\beta$-lactamases, 113 for cytochrome c, 238 for GFP, and 170 for Ras proteins.
-3.3 Backbone Dihedral Angles
+However, many methods generate structures that appear novel, diverse, and designable, yet few assess their functionality or evolutionary relevance, leaving the observations potentially as mere (Ji et al., 2023).
+Unclear biological functionality, unknown stability, and the lack of validation connecting these structures to known functions limit the broader application and advancement of these methods.
+We tackle these challenges using a deep generative pipeline, based on diffusion-based Score Matching (SM) and Flow Matching (FM), for early de novo design of monomeric proteins.
+Guided by an adaptable evaluation protocol, we apply it to four protein families chosen for their diverse structural folds, functional roles, and rich annotations (Appendix C).
+Generated structures are realistic and clash-free, and structural phylogenetic analyses show that they capture family-specific features consistent with common ancestry and function.
+Designed sequences conserve key functional residues while allowing variability elsewhere, yielding low-similarity variants with similar functions.
+Molecular Dynamics (MD) simulations demonstrate dynamic stability of the designs, while forming wild-type (WT)-like binding pockets that favorably interact with family-specific ligands in docking studies.
+Together, these results offer practical guidelines for integrating generative models into protein design workflows.
+2 Pipeline and Evaluation Protocol
+2.1 Protein Backbone Generation
+In Appendix B, we review the key concepts behind the approaches of Yim et al. (2023) and Bose et al. (2023) for backbone generation, using SM and FM on $\mathrm{SE}(3)$.
+2.2 Geometric Plausibility
 Due to steric hindrance and spatial repulsion, not all backbone dihedral angles in proteins are physically feasible or energetically favorable.
-We analyzed the $\phi$ and $\psi$ distribution in the generated structures using Ramachandran plots (Ramachandran et al., 1963) (Figure 1B).
-Most data points fall within the allowed and favored regions, with only a few in the disallowed areas, and no significant geometric clashes or unreasonable conformations were observed.
-The dihedral angle distributions of the generated structures align well with those of the experimental training proteins.
-For instance, cytochrome c structures have sparse points in the region $-180^{\circ}<\psi<-90^{\circ}$ and $45^{\circ}<\phi<180^{\circ}$, while GFP structures form four clusters there.
-SM samples are closer to the training data and concentrate in allowed regions but show less diversity than FM samples.
+We analyzed the $\phi$ and $\psi$ distribution in the generated structures using Ramachandran plots (Ramachandran et al., 1963) (Figure 6).
 2.3 Conserved Residue Consistency
-Figure 2: (A) Normalized evolutionary rates mapped onto structures, with black for rapidly evolving positions and orange for conserved ones.
-Experimental structures are highlighted with bold outlines, and highly conserved residues are marked in red (exposed) and green (buried).
-(B) Summary tree of two structural phylogenetic trees (Figure 7), constructed using $Q_{\text{score}}$ and the 3Di alphabet.
-Protein family members are grouped in semi-circular rings, with specific subtypes labeled in white.
-In $beta$-lactamases, different Ambler classes are shown with colored backgrounds.
-Colored branches and nodes indicate generated protein structures.
-In protein families or across species, certain residues are highly conserved, typically to ensure structural stability and support proper folding and function.
+In protein families or across species, certain residues are highly conserved, typically to ensure function but also structural stability and to support proper folding.
 Determining the optimal amino acid sequence.
@@
 (2023), we used ProteinMPNN (Dauparas et al., 2022) to predict ten sequences for each generated backbone.
-We then modeled these sequences with EMSfold (Rives et al., 2019) and compared them to the original backbone using the $\text{TM}_{\text{score}}$ (Zhang, 2005).
+We then modeled these sequences with EMSFold (Rives et al., 2019) and compared them to the generated backbone using the $\text{TM}_{\text{score}}$ (Zhang, 2005).
 The sequence with the highest $\text{TM}_{\text{score}}$ was chosen as the optimal match.
 Identifying conserved residues.
-The experimental protein sequences were aligned with each generated backbone’s optimal sequence using Clustal Omega (Sievers et al., 2011).
-Given these multiple sequence alignments (MSAs) and generated structures, ConSurf (Yariv et al., 2023) reconstructed phylogenetic trees and applied Rate4Site (Pupko et al., 2002) to calculate evolutionary rates for each position with an empirical Bayesian method (Ashkenazy et al., 2016; Mayrose, 2004).
-The optimal predicted sequences for the generated backbones showed conserved residues matching those in the experimental sequences (Figure 8).
-FM samples displayed greater diversity than SM samples, suggesting higher novelty, as noted in Section 3.5.
-Interestingly, this trend reverses in generated $\beta$-lactamases, possibly due to the greater structural variability in the experimental $\beta$-lactamase training set (Figure 2B), to which SM models are more sensitive.
-Section 3.6 further assesses the dynamic stability and ligand-binding properties of these generated samples.
-In Figure 2A, similar to the experimental proteins, the generated structures have higher residue conservation near binding pockets.
-For instance, in cytochrome c (1HRC) and generated structures SM-1 and FM-0, conserved residues cluster around the central heme C binding site.
-A similar pattern is seen in 4OBE and the generated KRas proteins.
+Experimentally derived sequences were aligned with optimal sequences of generated backbones using Clustal Omega (Sievers et al., 2011).
+Given these alignments and generated structures, ConSurf (Yariv et al., 2023) reconstructed phylogenetic trees and applied Rate4Site (Pupko et al., 2002) to estimate per-residue evolutionary rates via an empirical Bayesian method (Ashkenazy et al., 2016; Mayrose, 2004).
 2.4 Structural Phylogenetics
-In specific tasks, we aim to generate structures with dissimilar sequences while preserving similar functions.
-However, when sequence similarity falls below 30%, detecting homology and evolutionary relationships becomes challenging (Puente-Lelievre et al., 2023).
-Structural comparison, being more conserved (Illergård et al., 2009), is more effective for uncovering evolutionary connections (Flores et al., 1993; Puente-Lelievre et al., 2023; Moi et al., 2023).
+Certain applications require designing structures with low sequence similarity while retaining similar functions.
+However, when sequence identity falls below 30%, homology detection and evolutionary inference become challenging (Puente-Lelievre et al., 2023).
+Structural comparisons, which are more conserved, offer a more effective alternative (Illergård et al., 2009; Flores et al., 1993; Moi et al., 2023).
 Using $Q_{\text{score}}$ in structural phylogenetics.
 Malik et al.
-(2020) proposed the $Q_{\text{score}}$ (Krissinel & Henrick, 2004) for structural phylogenetics, as it accommodates indels and combines alignment quality with length.
-$Q_{\text{score}}$ compares the positions of all $\mathrm{C}_{\alpha}$ atoms among $N_{\text{align}}$ comparable residues in pairwise comparisons.
-We construct structural phylogenetic trees using $1-Q_{\text{score}}$ as a distance measure, with higher values indicating greater structural similarity.
-For any two structures with $N_{1}$ and $N_{2}$ residues, we align and calculate $Q_{\text{score}}$ using TM-align (Zhang, 2005) as:
-$Q_{\text{score}}=\frac{N_{\text{align}}^{2}}{N_{1}N_{2}}\times\frac{1}{1+\left(\frac{\text{RMSD}}{R_{0}}\right)^{2}}$ (1)
-[…]
+(2020) proposed the $Q_{\text{score}}$ (Krissinel & Henrick, 2004) for structural phylogenetics, which accommodates indels and combines alignment quality with length.
+It compares the positions of all $\mathrm{C}_{\alpha}$ atoms across $N_{\text{align}}$ comparable residues in pairwise comparisons.
+We construct structural phylogenetic trees using $1-Q_{\text{score}}$ as a distance measure, where higher values indicate greater structural similarity.
 Using 3Di alphabet in structural phylogenetics.
 van Kempen et al.
-(2023) developed Foldseek, which represents protein tertiary interactions as sequences over a 20-state 3D interaction (3Di) alphabet, simplifying structural alignments.
-This 3Di alphabet reduces false positives and increases information density by encoding conserved core regions more effectively.
-Leveraging Foldseek’s structural divergence metrics, Moi et al.
+(2023) developed Foldseek, which encodes protein tertiary interactions using a 20-state 3D interaction (3Di) alphabet to simplify structural alignments.
+This approach reduces false positives and increases information density by effectively encoding conserved core regions.
+Leveraging Foldseek’s divergence metrics, Moi et al.
 (2023) constructed structural phylogenetic trees based on rigid body alignment, local alignment without superposition, and sequence alignment with structural alphabets, showing that these trees outperform traditional sequence-based trees across varied evolutionary timescales (Moi et al., 2023).
-Summary tree.
-Two methods were used to calculate pairwise structural distances between generated and experimental structures, creating a distance matrix for constructing structural phylogenetic trees (Figure 7).
-After normalizing branch lengths, we generated a summary tree using SumTrees (Moreno et al., 2024), shown in Figure 2B.
-To ensure reliable $Q_{\text{score}}$ calculations, residue count differences between compared structures were limited to within 10% (Malik et al., 2020).
-Thus, we included proteins with residue counts of 243-297 for $\beta$-lactamases, 102-124 for cytochrome c, 216-261 for GFP, and 166-187 for Ras proteins, excluding class C $\beta$-lactamases ($\sim 360$ residues) from the tree but retaining them for model training and analysis.
-The summary tree illustrates the potential of structural phylogenies, with each protein family clustering on distinct branches and subtypes (such as metallo-$\beta$-lactamases) grouping closely.
-Generated structures mostly integrate within small branches of experimental proteins, suggesting a close structural relationship.
-However, the models may not fully capture the specificity of class D $\beta$-lactamases, as no generated samples clustered in this class.
-SM samples form dense clusters, while FM samples show greater diversity, positioning closer to the tree root but near experimental proteins without creating new branches, indicating higher novelty in FM outputs.
 2.5 Molecular Dynamics
-Figure 3: Stability assessment of MD simulations across proteins using various metrics.
-(A, C-E) Distributions of (A) RMSD, (C) radius of gyration ($R_{g}$), (D) potential energy, and (E) secondary structure counts throughout the simulation.
-Interquartile ranges and whiskers show metric variation; high-quality structures have medians close to experimental values with narrow ranges.
-(B) Residue fluctuation (RMSF) during simulation, where stable structures exhibit lower RMSF values and trends similar to the experimental structure.
-Structures that appear reasonable may, in fact, be unstable due to molecular dynamics, water interactions, and entropy, which are not considered during generation.
-To evaluate the dynamic stability of the generated structures, we conducted MD simulations under physiological conditions and analyzed their time-dependent behavior.
+Structures that appear reasonable may, in fact, be unstable due to molecular dynamics, water interactions, and entropy, which are not accounted for during generation.
+To assess the dynamic stability of these generated structures, we conducted MD simulations under physiological conditions and analyzed their time-dependent behavior.
 Homology modeling of side-chains.
@@
 We chose the final side-chain arrangement based on lowest energy and minimal steric clashes, verifying model quality with PROCHECK (Laskowski et al., 1993, 1996) and WHAT_CHECK (Hooft et al., 1996), discarding low-quality models.
+MODELLER’s modeling leverages (a) the input alignment to position side-chains, (b) template structures to set spatial restraints that mimic contacts, (c) knowledge-based rotamer distributions to favor typical angles, and (d) an optimization process.
 Simulation.
-After adding hydrogen atoms with Reduce2 from the computational crystallography toolbox (Grosse-Kunstleve et al., 2002), we conducted MD simulations with GROMACS (Abraham et al., 2015).
-The protein was placed in an octahedral box with at least 15 Å̊$\mathring{\mathrm{A}}$ from the edges, using the CHARMM36 force field (July 2022) (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Yu et al., 2012; Soteras Gutiérrez et al., 2016) for intermolecular interactions.
-Following vacuum energy minimization, the system was solvated, neutralized to 150 mM Na+ and Cl-, and minimized again.
-The system was then heated to 310 K under NVT [^2] Constant number of particles, volume, and temperature.
-conditions and equilibrated at 1 atm under NPT [^3] Constant number of particles, pressure, and temperature.
-conditions, followed by a 10 ns production run.
-Details are provided in Appendix F.
-Stability.
-Dynamic stability were analyzed as follows:
+Hydrogen atoms were added using Reduce2 from the computational crystallography toolbox (Grosse-Kunstleve et al., 2002), and MD simulations were performed with GROMACS (Abraham et al., 2015).
+Proteins were placed in an octahedral box with at least 15 Å̊$\mathring{\mathrm{A}}$ from the edges, and intermolecular interactions were modeled using the CHARMM36 force field (July 2022) (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Yu et al., 2012).
+After vacuum energy minimization, the system was solvated, neutralized to 150 mM Na+ and Cl-, and minimized again.
+Then it was heated to 310 K under NVT [^1] Constant number of particles, volume, and temperature., equilibrated at 1 atm under NPT222constant number of particles, pressure, and temperature, and subjected to a 10 ns production run.
+Details are provided in Appendix I.
+2.6 Protein-ligand Docking
+AutoDock Vina (Trott & Olson, 2009; Eberhardt et al., 2021) was used to predict optimal binding modes between generated structures and their family-specific ligands.
+We performed blind docking, scanning the entire protein surface for potential ligand binding sites without prior knowledge of the pockets.
+The grid box covered the entire protein, and identical configurations were applied to both generated and experimentally derived structures to evaluate whether deviations in the generated samples fell within an acceptable range, thus assessing their functional viability.
+Details on simulation settings are provided in Section I.3.
+3 Case Studies
+3.1 Data
+We assembled a dataset of monomeric proteins covering four families ($\beta$-lactamases, cytochrome c, GFP, and Ras), with varied fold types and functions, incorporating both natural and engineered mutations while retaining conserved core functional regions.
+Details are provided in Appendix C.
+3.2 Backbone Generation
+We trained the SM and FM (with Optimal Transport) models on these protein families, using pretrained weights from Yim et al.
+(2023) and Bose et al.
+(2023), with each model having $\sim 17$ million parameters.
+Each model generated 50 backbone structures, with target sequence lengths sampled from the distribution observed in the training data (Figure 5).
+3.3 Backbone Dihedral Angles
+In Figure 6, most data points fall within the allowed and favored regions, with few in the disallowed areas, and no significant geometric clashes or unreasonable conformations observed.
+The dihedral angle distributions of the generated structures align well with those of the experimentally derived proteins used for training.
+For instance, cytochrome c structures have sparse points in the region $-180^{\circ}<\psi<-90^{\circ}$ and $45^{\circ}<\phi<180^{\circ}$, while GFP structures form four clusters there.
+SM samples are closer to the training data and concentrate in allowed regions and show less diversity than FM samples.
+3.4 Conserved Residue Consistency
+The optimal sequences preserved conserved residues that were largely consistent with the experimentally derived data (Figure 8).
+In Figure 7, similar to the experimentally derived proteins, the generated structures have increased residue conservation around binding pockets (refer to Figure 4 and Figure 12).
+For instance, in cytochrome c (1HRC) and generated SM-1 and FM-0, conserved residues cluster near the central heme C binding site.
+A similar pattern appears in 4OBE and the generated KRas proteins.
+3.5 Structural Phylogenetic Tree
+Summary tree.
+Pairwise structural distances between generated and experimental structures were computed using two methods, forming a matrix for phylogenetic tree construction.
+After normalizing branch lengths, a summary tree was generated with SumTrees (Moreno et al., 2024) (Figure 9).
+The topology shows that both the generated and the experimentally derived proteins cluster strongly by family, without intermixing among these groups.
+Structure-informed tree outperforms sequence-only tree.
+Sequence-based phylogenetic trees were constructed by aligning sequences with Clustal Omega and inferring trees with FastTree (Price et al., 2009).
+Structure trees better preserve natural taxonomic groupings than sequence trees (Figure 10AB), especially at moderate to low sequence identity, and they achieve higher Taxonomic Congruence Scores (Tan et al., 2015) (Appendix G, Figure 10D), indicating closer alignment with the known taxonomy (Appendix G).
+3.6 Dynamic Stability
 […]
 3.7 Protein-ligand Docking
-Figure 4: Comparison of experimentally determined bindings (bold boxes) with predictions from blind docking simulations, using models closest to experimental data.
-Ligands are shown in green (experimental data), cyan (predictions from experimental receptors), magenta (predictions from SM samples), and yellow (predictions from FM samples).
-$\Delta G$ is in kcal/mol, with RMSD indicating deviation from experimental ligand positions (in Å).
-Dockings labeled “MD” represent protein-ligand complex MD simulations after blind docking.
-Using AutoDock Vina (Trott & Olson, 2009; Eberhardt et al., 2021), we predicted optimal binding modes between generated structures and their family-specific ligands.
-We performed blind docking, scanning the entire protein surface for potential ligand binding sites without prior knowledge of binding pockets.
-The grid box covered the entire protein, and identical configurations were applied to both generated and experimental structures to evaluate if deviations in generated samples fell within an acceptable range, thus assessing their functional viability.
-Details on simulation settings are provided in Appendix G.
-[…]
 […]
 $\beta$-lactamases binding to penicillin.
-Mutations at Glu166 and Asn170 in class A $\beta$-lactamases (Figure 1C) can form a stable acyl-enzyme intermediate, disrupting deacylation (Chen & Herzberg, 2001).
-Only asparagine and glycine at position 170 preserve WT-like function (Brown et al., 2009), and this conservation is retained in generated structures like FM-4 (Figure 4A), suggesting they may retain the ability to inactivate penicillin antibiotics.
+Mutations at Glu166 and Asn170 in class A $\beta$-lactamases (Figure 4A) can form a stable acyl-enzyme intermediate, disrupting deacylation (Chen & Herzberg, 2001).
+Only Asn and Gly at 170 preserve WT-like function (Brown et al., 2009), and this conservation is retained in generated samples like FM-4 (Figure 12A), suggesting they may retain the ability to inactivate penicillin antibiotics.
+Notably, Asn/Gly at position 170 occurred in 23/50 FM samples and 18/50 SM samples, and when extended to positions 169 to 171, in 30/50 FM samples and 22/50 SM samples, rates higher than expected by chance.
 Cytrochrome c binding to heme c.
-The generated cytochrome c-like structures retain conserved residues found in the WT (Figure 1D; 1HRC).
+The generated cytochrome c-like structures retain conserved residues found in the WT (Figure 4B; 1HRC).
 SM-1 (Figure 12B) includes phosphorylatable residues Tyr58, Thr59, and Tyr107, as well as lysine residues Lys82, Lys83, and Lys96 involved in phospholipid binding.
@@
 GNP, a non-hydrolyzable GTP analog, is commonly used to simulate the GTP-bound active state of Ras proteins (Pantsar, 2020).
-In WT KRas bound to GNP (Figure 4D; 5UFE), residues from the Switch II region fill the pocket.
-Removal of the $\gamma$-phosphate relaxes the protein conformation to the GDP-bound state, opening the switch II region (Figure 4C; 4OBE) (Kauke et al., 2017).
+In WT KRas bound to GNP (Figure 12D; 5UFE), the Switch II region residues fill the pocket.
+Removing of the $\gamma$-phosphate relaxes the conformation to the GDP-bound state, opening the switch II region (Figure 12C; 4OBE) (Kauke et al., 2017).
 After blind docking the generated KRas-like structures with GNP/GDP, we analyzed the complexes using protein–ligand MD simulations.
-Allowing flexibility in both the backbone and side-chains enabled us to capture the dynamic conformational changes, especially in the switch regions, between active/inactive states.
-FM-generated KRas-like structures show greater flexibility.
+Allowing flexibility in both the backbone and side-chains enabled us to capture the dynamic conformational changes, especially the WT-like transitions in the switch regions, between active/inactive states.
+FM-generated structures show greater flexibility.
 In FM-13 (Figure 12D), GNP binding causes switch II residues to fill the pocket, similar to 5UFE.
@@
 4 Discussion
-This study demonstrates the potential of SM and FM methods for generating protein backbones.
-After determining the sequence, we performed side-chain homology modeling.
-Despite sequence dissimilarities, key residues are conserved, and the generated structures are highly similar.
-SM better captures conserved regions, producing more rigid structures, while FM offers greater flexibility.
-Structural phylogenetic analysis revealed evolutionary relationships, and MD and docking simulations confirmed stability and functionality.
-The presented pipeline and evaluation protocols build confidence for broader applications in protein generation.
 SM and FM can generate a range of monomeric protein structures and can have applications beyond protein design (Yim et al., 2023; Bose et al., 2023).
@@
 Although some progress has been made (Torge et al., 2023; Jin et al., 2023; Somnath et al., 2023; Zhou et al., 2023), research gaps remain.
-These generative processes differ from protein synthesis simulations and do not reflect human hierarchical understanding.
-Limited interpretability reduces their utility for further downstream analyses and industrial applications.
-Although generated proteins may resemble real ones and include similar conserved residues and conformational changes, small variations could significantly impact function.
-Simply aiming for diversity and novelty does not guarantee quality.
-Generative models have also shown great potential as efficient surrogates for MD simulations, which are often computationally intensive for complex systems or long-term behaviors.
-Jing et al.
-(2024) developed a framework to simulate molecular trajectories using generative models.
-Similarly, Viguera Diez et al.
-(2023) used generative models to enhance sampling of slow degrees of freedom, covering the sample space more effectively than traditional MD.
+The generated samples for the most part capture the observed evolutionary diversity.
+However, there are regions of the observed diversity that are not so well covered (Figure 9) (such as GFP and class A $\beta$-lactamases), raising concerns about potential overfitting and limited generalizability.
+Common metrics are inadequate for quantifying overfitting in generative models (Arora et al., 2017), especially given the variability in sample sizes and quality across families.
+5 Guidelines
+Here, we outline a set of guidelines and best practices for designing and validating proteins using deep generative methods: (1) Implement physical constraints during the generation process to produce realistic structures without severe steric clashes.
+(2) Generate multiple candidate sequences for each design and prioritize those most likely to fold into the target structure.
+(3) Integrate sequence design and backbone validation early to ensure compatibility between the sequence and the backbone model.
+(4) Identify and preserve conserved residues in the designs to maintain critical functional motifs and overall structural integrity.
+(5) Construct all-atom models of the designs and refine them to relieve any strain or clashes introduced during initial modeling.
+(6) Compare the designs to known protein family structures or fold templates to ensure consistency with known protein folds.
+(7) Perform molecular dynamics simulations to verify that the proposed structures remain stable under physiologically relevant conditions.
+(8) Perform docking analyses to evaluate the viability of proposed ligand-binding sites.
+(9) Experimental validation is always essential to confirm the intended function and structural integrity of each candidate.
+(10) See Appendix J for circumstances under which implementation of these guidelines can be challenging.
+6 Conclusion
+This study shows the potential of methods based on deep generative models for designing proteins with high structural fidelity and functional plausibility.
+SM better captures conserved regions, producing more rigid structures, while FM offers greater flexibility (Figure 8).
+Although the optimized sequences for the generated structures exhibit some differences from natural family members, they preserve many functionally essential conserved residues (Figure 7).
+In structural phylogenetic trees (Figure 9), these designs cluster with their respective families, suggesting that their geometries could capture some family‐specific signatures indicative of common ancestry or function.
+MD and docking simulations are consistent with the designs remain properly folded and stable under physiological conditions over time (Figure 11), and can accommodate their family-specific ligands in binding pockets that closely correspond to known active sites (Figure 12).
+Some designs undergo conformational changes upon binding that mirror the allosteric behavior observed in WT (Figure 12CD).
 Software and Data
@@
 The PDB structure files were downloaded on June 19, 2024, from https://www.wwpdb.org/ftp/pdb-ftp-sites.
-Acknowledgments
-The authors would like to thank Charlie Harris and Gos Micklem for their fruitful discussions and valuable feedback on this work.
-Appendix A Protein Backbone Generation
-A.1 $\mathrm{SE}(3)$ Decomposition into $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$
-The Special Euclidean group $\mathrm{SE}(3)$ describes the rotations and translations in 3D space. An element of the $\mathrm{SE}(3)$ can be represented by a $4\times 4$ matrix:
+Acknowledgements
+The authors would like to thank Charlie Harris, Simon Mathis and Stephen Eglen for their fruitful discussions and valuable feedback on this work.
+Appendix A Deep Generative Protein Design Workflow
+Figure 1: A schematic overview of the deep generative protein design pipeline and its evaluation protocols.
+Appendix B Protein Backbone Generation
+B.1 $\mathrm{SE}(3)$ Decomposition into $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$
+The Special Euclidean group $\mathrm{SE}(3)$ describes the rotations and translations in 3D space. An element of $\mathrm{SE}(3)$ can be represented by a $4\times 4$ matrix:
 $\bm{\mathsf{T}}=\begin{pmatrix}\bm{\mathsf{R}}&\bm{\mathsf{x}}\\ \mathbf{0}_{1\times 3}&1\end{pmatrix}$ (1)
@@
 Since $\mathrm{SE}(3)$ can be viewed as the semidirect product of $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$, denoted as $\mathrm{SE}(3)\cong\mathrm{SO}(3)\ltimes\mathbb{R}^{3}$, one option is to naturally treat $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$ as independent for simplicity (Yim et al., 2023).
-A.2 Protein Backbone Representations
+B.2 Protein Backbone Representations
+Figure 2: Protein backbone with dihedral angles $\psi$ and $\phi$.
 Molecules can be intuitively represented as 3D atomic point clouds.
@@
 (2023), we adopt the more compact backbone rigid groups from AlphaFold (Jumper et al., 2021) to represent protein backbone structures in 3D space.
-A backbone rigid group consists of the main chain atoms (N, Cα, C, O) within a single residue (Figure 1A), where the their geometric relationships (relative positions and orientations) are highly stable.
+A backbone rigid group consists of the main chain atoms (N, Cα, C, O) within a single residue (Figure 2), where their geometric relationships (relative positions and orientations) are highly stable.
 The position and orientation of the group is transformed as a whole, without accounting for individual atomic movements, simplifying the computation and reducing structural errors caused by excessive degrees of freedom.
-Assuming experimentally measured ideal chemical bond angles and lengths, models learn how the rigid transformation (or frame) $\bm{\mathsf{T}}_{i}$ of each residue $i\in[1,N]$ acts on idealized coordinates $[\mathrm{N}^{\star},\mathrm{C}_{\alpha}^{\star},\mathrm{C}^{\star}]\in\mathbb{R}^{3}$ (centered at $\mathrm{C}_{\alpha}^{\star}$), so that the transformed coordinates match the actual coordinates as closely as possible:
+Assuming experimentally derived ideal chemical bond angles and lengths, models learn how the rigid transformation (or frame) $\bm{\mathsf{T}}_{i}$ of each residue $i\in[1,N]$ acts on idealized coordinates $[\mathrm{N}^{\star},\mathrm{C}_{\alpha}^{\star},\mathrm{C}^{\star}]\in\mathbb{R}^{3}$ (centered at $\mathrm{C}_{\alpha}^{\star}$), so that the transformed coordinates match the actual coordinates as closely as possible:
 $[\mathrm{N},\mathrm{C}_{\alpha},\mathrm{C}]_{i}=\bm{\mathsf{T}}_{i}\cdot[\mathrm{N}^{\star},\mathrm{C}_{\alpha}^{\star},\mathrm{C}^{\star}]$ (2)
@@
 An additional torsion angle $\psi_{i}\in\mathrm{SO}(2)$ is introduced between the bond of $\mathrm{C}_{\alpha}$ and C for a more accurate construction of the backbone oxygen atom O.
-A.3 $\mathrm{SE}(3)$ Score Matching
+B.3 $\mathrm{SE}(3)$ Score Matching
 Let $\mathbf{T}_{t}=\left[\bm{\mathsf{T}}_{1,t},\dots,\bm{\mathsf{T}}_{N,t}\right]\in\mathrm{SE}(3)^{N}$ denote the manifold of $N$ frames at time $t$, where each frame can independently rotate and translate.
@@
 However, this gradient is typically intractable in practice because the exact form of $p_{t}(\mathbf{T}_{t})$ is unknown at any given time $t$.
-Instead, score-based models estimates tractable conditional score $\nabla\log p_{t|0}$ through SM (Vincent, 2011), using a neural network $s(\theta,t,\cdot)$ trained by minimizing both [^4] One adds weights $1/\mathbb{E}[\lVert\nabla_{\mathbf{R}}\log p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})\rVert^{2}]$ to Equation 7 and $(1-e^{-t})/e^{-t/2}$ to Equation 8 for simplicity.:
+Instead, score-based models estimate tractable conditional score $\nabla\log p_{t|0}$ through SM (Vincent, 2011), using a neural network $s(\theta,t,\cdot)$ trained by minimizing both [^3] One adds weights $1/\mathbb{E}[\lVert\nabla_{\mathbf{R}}\log p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})\rVert^{2}]$ to Equation 6 and $(1-e^{-t})/e^{-t/2}$ to Equation 7 for simplicity.:
 $\mathcal{L}_{\text{SM}}^{\mathbf{R}}(\theta)$ $=\mathbb{E}\left[\lVert\nabla_{\mathbf{R}}\log p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})-s(\theta,t,\mathbf{R}_{t})\rVert^{2}\right]$ (6)
@@
 of the Brownian motion on $\mathrm{SO}(3)$.
-A.4 $\mathrm{SE}(3)$ Flow Matching
+B.4 $\mathrm{SE}(3)$ Flow Matching
 FM is a simulation-free method for training vector fields to follow a prescribed conditional probability path (Lipman et al., 2022).
@@
 $\mathcal{L}_{\text{OT}}^{\mathbf{R}}(\theta)=\mathbb{E}_{\bar{\pi}}\left\lVert\frac{\log_{\mathbf{R}_{t}}(\mathbf{R}_{0})}{t}-v(\theta,t,\mathbf{R}_{t})\right\rVert^{2}$ (18)
-A.5 $\mathrm{SE}(3)$ Invariance
+B.5 $\mathrm{SE}(3)$ Invariance
 $\mathrm{SE}(3)$ invariance can be achieved by consistently positioning the model at the origin (Yim et al., 2023; Bose et al., 2023; Rudolph et al., 2020).
-In the context of SM, to ensure translation invariance on $\mathbb{R}^{3}$, one apply a projection matrix $\mathbf{P}\in\mathbb{R}^{3N\times 3N}$ that removes the center of mass $\frac{1}{N}\sum_{i=1}^{N}\bm{\mathsf{x}}_{i}$.
+In the context of SM, to ensure translation invariance on $\mathbb{R}^{3}$, one applies a projection matrix $\mathbf{P}\in\mathbb{R}^{3N\times 3N}$ that removes the center of mass $\frac{1}{N}\sum_{i=1}^{N}\bm{\mathsf{x}}_{i}$.
 It results in an invariant measure on $\mathrm{SE}(3)^{N}$, denoted as $\mathrm{SE}(3)^{N}_{0}$.
@@
 After centering and decoupling the flow on $\mathrm{SE}(3)_{0}^{N}$, a separate $\mathrm{SE}(3)$-invariant flow can be constructed for each residue in backbone [^5] As the product group of $N$ copies of $\mathrm{SE}(3)$, $\mathrm{SE}(3)_{0}^{N}$ has a geometric structure that allows global geometric operations (such as geodesic distance, exponential maps, and logarithmic maps) to be decomposed into operations on each of the $N$ $\mathrm{SE}(3)$ groups., in which each $\mathrm{SE}(3)$-invariant measure is decomposed into a measure that is proportional to the Lebesgue measure on $\mathbb{R}^{3}$ (Pollard, 2001) and an $\mathrm{SO}(3)$-invariant measure (Bose et al., 2023).
-A.6 Additional Losses
+B.6 Additional Losses
 To prevent unrealistic fine-grained features such as steric clashes or chain breaks when learning the torsion angle $\psi$, Yim et al.
@@
 where $T=1$ in the case of FM.
-A.7 Model Architecture
+B.7 Model Architecture
 Figure 3: Overview of the (A) embedding module and (B) multi-layer network architecture.
-The networks $s(\theta,t,\cdot)$ involved in SM and $v(\theta,t,\cdot)$ involved in FM models, as reviewed in Section 2, can share a common high-level architecture (Yim et al., 2023; Bose et al., 2023; Anand & Achim, 2022).
+The networks $s(\theta,t,\cdot)$ involved in SM and $v(\theta,t,\cdot)$ involved in FM models, as reviewed in Appendix B, can share a common high-level architecture (Yim et al., 2023; Bose et al., 2023; Anand & Achim, 2022).
 Embeddings.
-Given node embedding dimensions $D_{h}$ and edge embedding dimensions $D_{e}$, node embeddings $\mathbf{h}\in\mathbb{R}^{N\times D_{h}}$ are derived from residue indices $\textbf{i}=\{1,\dots,N\}$ and time-step information $\textbf{t}=\{0,\Delta t,\dots,T\}$, while edge embeddings $\mathbf{E}=\in\mathbb{R}^{N\times N\times D_{e}}$ integrate additional features, such as relative sequence distances $j-i$ for any $i,j\in[1,N]$ (Figure 5A).
+Given node embedding dimensions $D_{h}$ and edge embedding dimensions $D_{e}$, node embeddings $\mathbf{h}\in\mathbb{R}^{N\times D_{h}}$ are derived from residue indices $\textbf{i}=\{1,\dots,N\}$ and time-step information $\textbf{t}=\{0,\Delta t,\dots,T\}$, while edge embeddings $\mathbf{E}\in\mathbb{R}^{N\times N\times D_{e}}$ integrate additional features, such as relative sequence distances $j-i$ for any $i,j\in[1,N]$ (Figure 3A).
 Self-conditioning on the predicted $\mathrm{C}_{\alpha}$ displacements is also applied:
 $c_{i,j}=\sum_{b=1}^{B}\mathds{1}\left\{|\bm{\mathsf{x}}_{i}^{*}-\bm{\mathsf{x}}_{j}^{*}|<v_{b}\right\}$ (25)
-where $\bm{\mathsf{x}}^{*}$ denotes the coordinates for $\mathrm{C}_{\alpha}$ predicted through self-conditioning, and $v_{1},\dots,v_{B}$ are bins spaced uniformly from 0 and $B$ angstroms.
+where $\bm{\mathsf{x}}^{*}$ denotes the coordinates for $\mathrm{C}_{\alpha}$ predicted through self-conditioning, and $v_{1},\dots,v_{B}$ are bins spaced uniformly from 0 to $B$ angstroms.
 These initial features are encoded using multilayer perceptrons (MLPs) along with sinusoidal embeddings (Vaswani et al., 2017).
@@
 Torsion angle $\bm{\hat{\psi}}=\{\hat{\psi}_{1},\dots,\hat{\psi}_{N}\}=\bm{\psi}/\lVert\bm{\psi}\rVert\in\mathrm{SO(2)}^{N}$ is predicted with $\textbf{h}^{L}$ and $\textbf{E}^{L}$.
-Appendix B Amino Acid Count Distributions in Experimental Proteins Used for Training
-Figure 6: (A-D) Distribution of amino acid sequence lengths (aa) in experimentally determined protein structures used for training.
-Appendix C Structural Phylogenetic Trees
-Figure 7 shows structural phylogenetic trees generated using $Q_{\text{score}}$ and the 3Di alphabet.
-Both approaches reveal evolutionary relationships between the generated and experimental structures.
-However, compared to the $Q_{\text{score}}$-based tree, the 3Di alphabet tree shows closer and less distinguishable evolutionary relationships among samples from different protein families, especially between the GFP and cytochrome c samples.
-Figure 7: Structural phylogenetic trees constructed using (A) $Q_{\text{score}}$ and (B) 3Di alphabet.
-Members of different protein families are grouped into semi-circular rings.
-The colored branches and nodes represent the generated protein structures.
-Appendix D Conserved Residue Consistency
-Figure 8: Normalized evolutionary rates were mapped onto sequences, with black indicating rapidly evolving positions and orange indicating conserved positions.
-The heat map displays each column as the optimal predicted sequence of a generated sample, with rows representing specific positions in these sequences.
-The bar chart (exp.
-identity) shows the average pairwise sequence identity at each position from the MSAs of experimental proteins.
-Taller bars signify positions where most experimental protein share the same residue.
-Consistency between generated and experimental sequences at conserved positions appears as uniform orange coloring in the heat map, aligning with the taller bars.
-Appendix E Evaluation of Side-Chain Homology Modeling
+Appendix C Protein Families Involved in This Study
+Figure 4: (A) Overlay of WT E.
+coli TEM1 (PDB: 1BTL; Jelsch et al.
+(1993); green) and its E166N acylated intermediate (PDB: 1FQG; Brown et al.
+(2009); white) with penicillin (PNM).
+(B) WT E.
+caballus heart cytochrome c (PDB: 1HRC; (Dickerson et al., 1967)) with heme C (HEC).
+(C) A.
+victoria GFP and its chromophore (PDB: 4KW4; Barnard et al.
+(2014)).
+(D) GDP-bound H.
+sapiens KRas protein (PDB: 4OBE; (Hunter et al., 2014)).
+C.1 $\beta$-lactamases
+$\beta$-lactamases are enzymes that deactivate $\beta$-lactam antibiotics by hydrolyzing their $\beta$-lactam ring, contributing significantly to bacterial resistance (Lee et al., 2016).
+Inhibiting these enzymes can restore antibiotic efficacy (Behzadi et al., 2020).
+Rapid diversification driven by the evolution of bacterial resistance makes $\beta$-lactamases ideal targets for protein modeling studies.
+For this study, we gathered structural data on 1,578 unique monomeric $\beta$-lactamases and their variants across Ambler classes (A, B, C, and D) from the BLDB (Naas et al., 2017) and the Protein Data Bank (PDB) (Berman, 2000).
+Class A $\beta$-lactamases are the most prevalent, with conserved active-site residues Ser70, Glu166, and Asn170 coordinating the hydrolytic water for deacylation (Tooke et al., 2019; Brown et al., 2009).
+Figure 4A shows the WT $\beta$-lactamase TEM1 (1BTL) alongside its acylated E166N intermediate (1FQG).
+C.2 Cytochrome c
+Cytochrome c is a water-soluble protein ($\sim 12$ kDa) essential for ATP synthesis in mitochondria and intrinsic apoptosis (Kashyap et al., 2021; Ow et al., 2008).
+It also serves as an independent marker for apoptosis in several cancers (Li et al., 2001; Way et al., 2004).
+Despite variations across species, its core structure and function are conserved.
+We obtained structural data for 498 unique cytochrome c proteins and variants from the PDB.
+Figure 4B shows horse cytochrome c, where a hydrophobic shell surrounds the heme group, with only $\sim 7.5\%$ of the surface available for electron transfer (Bushnell et al., 1990).
+The hydrophobic environment and iron coordination by His18 and Met80 maintain a high redox potential ( 260 mV) (Salemme, 1977).
+Phosphorylation occurs at Thr28, Thr47, Tyr48, and Tyr97 (Hüttemann et al., 2011), while Lys72, Lys73, and Lys87 bind phospholipids (Kagan et al., 2009).
+The ATP-binding pocket involves Glu69, Asn70, Lys88, and Lys72, Lys86, Lys87 (McIntosh et al., 1996).
+C.3 Green fluorescent proteins (GFP)
+GFP, first isolated from Aequorea victoria, fluoresces green when stimulated by specific wavelengths.
+Its core structure is an 11-strand $\beta$-barrel enclosing the chromophore (Figure 4C) (Remington, 2011).
+Various mutants have been engineered to enhance or modify its properties, including enhanced GFP (Cormack et al., 1996), superfolder GFP (Pédelacq et al., 2005), and color variants like YFP (Ormö et al., 1996) and BFP (Glaser et al., 2016).
+We collected structural data for 448 GFPs and variants from the PDB to explore this diversity.
+C.4 Ras Proteins
+Ras proteins, a subgroup of the small GTPase superfamily, act as molecular switches, cycling between GTP-bound (active) and GDP-bound (inactive) states to regulate cell proliferation, differentiation, migration, and apoptosis (Ladygina et al., 2011; Simanshu et al., 2017; Weinmann & Ottow, 2007).
+They play a key role in signaling from the cell surface to downstream pathways.
+Mutations that keep Ras proteins in an active state drive excessive cell growth and malignancy, making Ras inhibition a promising cancer treatment strategy (Singh & Lingham, 2002).
+We focused on the most common cancer-related Ras proteins (HRas, KRas, and NRas) (Cox, 2002) and obtained 511 experimentally derived structures from the PDB.
+In human KRas (Figure 4D), the switch I and II regions form the key interface for effector and regulator binding (Pantsar, 2020).
+These regions are highly flexible, with conformations depending on GTP or GDP binding.
+Cancer-related mutations frequently occur in the P-loop and switch II (Pantsar, 2020).
+C.5 Sequence Length Distributions in Experimentally Derived and Generated Protein Structures
+Figure 5: Distributions of amino acid sequence lengths (aa) for the experimentally derived protein structures used for training (top row; orange) and the 50 backbone structures generated by each model (bottom row; blue).
+Target sequence lengths for the generated structures were sampled from the training data distribution, shown here for (A) $\beta$-lactamase, (B) cytochrome c, (C) GFP, and (D) Ras.
+Appendix D Dihedral Angles $\psi$ and $\phi$ Distributions
+Figure 6: Ramachandran plots comparing dihedral angles $\psi$ and $\phi$ distributions for generated versus experimentally derived proteins; inset shows favored (light) and disallowed (dark) regions.
+Appendix E Conserved Residue Consistency
+E.1 Evolutionary Rates Mapped onto Structures
+Figure 7: Normalized evolutionary rates mapped onto structures, with white for rapidly evolving positions and black for conserved ones.
+Experimentally derived structures are highlighted with bold outlines.
+Refer to Figure 4 and Figure 12 for the relevant key residues and the locations of the ligand-binding pockets (represented by PNM, HMC and GDP).
+E.2 Evolutionary Rates Mapped onto Sequences
+Figure 8 shows that, except for $\beta$‑lactamase, the FM-generated sequences show a slightly higher average pairwise distance than those of the SM, indicating greater diversity.
+One possible explanation is that the experimentally derived $\beta$‑lactamase structures used for training (from four distinct Ambler classes) have more variability than the other three protein families (as also reflected in Figure 9) and that SM is more sensitive to structural variability.
+Figure 8: Multiple sequence alignment of generated samples and experimentally derived reference proteins (1FQG, 1HRC, 4XWA, 4O8E) using Clustal Omega, with normalized evolutionary rates overlaid.
+Rapidly evolving positions are highlighted in light gray, conserved regions in black, and alignment gaps in white.
+In addition, a comparative analysis of variability is presented: average pairwise distances, derived from the Clustal Omega distance matrix, denote greater variability when larger values are observed.
+Appendix F Summary Structural Phylogenetic Tree
+Figure 9: Summary structural phylogenetic tree constructed using the $Q_{\text{score}}$ and the 3Di alphabet.
+Branch colors distinguish families, with orange and blue nodes representing SM- and FM-generated structures, respectively.
+In $\beta$-lactamases, distinct Ambler classes are differentiated by unique background colors.
+Appendix G Structure-informed Trees versus Sequence-based Phylogenetic Trees
+Retrieving known taxonomic lineages.
+Following Moi et al.
+(2023), we retrieved taxonomic lineages for each experimentally derived sequence and structure within every protein family using the UniProt API (Patient et al., 2008), assuming that most genes evolve in a manner that mirrors the species tree with only occasional instances of gene loss or duplication.
+Figure 10: (A) Ultrametric summary structural phylogenetic tree constructed using the $Q_{\text{score}}$ and the 3Di alphabet.
+(B) Ultrametric sequence-based phylogenetic tree constructed using the Clustal Omega and the FastTree pipeline.
+Different protein families are differentiated by distinct colored branches.
+Nodes in orange and blue represent structures generated by SM and FM, respectively.
+(C) Phylogenetic trees of distinct protein families.
+Top: summary structural phylogenetic tree constructed using the $Q_{\text{score}}$ metric and the 3Di alphabet.
+Bottom: sequence-based phylogenetic tree inferred using the Clustal Omega and FastTree pipeline.
+(D) Taxonomic congruence score for each node in the sequence and structure trees.
+On average, structural trees exhibit higher taxonomic congruence than sequence-based trees.
+Using $Q_{\text{score}}$ in structural phylogenetics.
+For any two structures with $N_{1}$ and $N_{2}$ residues, $Q_{\text{score}}$ is computed with TM-align (Zhang, 2005) as:
+$Q_{\text{score}}=\frac{N_{\text{align}}^{2}}{N_{1}N_{2}}\times\frac{1}{1+\left(\frac{\text{RMSD}}{R_{0}}\right)^{2}}$ (34)
+[…]
+Taxonomic congruence score (TCS).
+Tan et al.
+(2015) proposed the use of the TCS to assess how well a phylogenetic tree’s topology agrees with the known taxonomy, arguing that TCS is an unbiased measure of tree quality.
+Typically, trees with higher average TCS (structure trees in Figure 10D) are considered to have more accurate topologies.
+In this study, we evaluated the congruence between the given trees and the established taxonomy lineages derived by UniProt.
+Subsequently, we provide a brief overview of the bottom-up TCS implementation as described by Moi et al.
+(2023):
+For any node $x$ in the tree, $s(x)$ is the set of taxonomic lineage labels present in the subtree rooted at $x$.
+If $x$ is a leaf, then $s(x)$ is defined to be the set of lineage labels (its taxonomic classification) for that leaf.
+If $x$ is an internal node with children $\{y_{1},\cdots,y_{N}\}$, then $s(x)=\bigcap_{i=1}^{N}s(i)$.
+For each node $x$, the function $C(x)$ quantifies the congruence of that node’s grouping with taxonomy: $C(x)=|s(x)|+|s(p)|$, where $p$ is the parent nodes of $x$, and ||$\cdot$|| ⋅ | denotes the size of a set.
+The overall taxonomic congruence score for the entire tree is obtained by summing the contributions of all leaves.
+To compare congruence across trees of different sizes, the raw total score is typically normalized.
+Appendix H Evaluation of Side-Chain Homology Modeling
 We used homology modeling to add side-chains to the generated protein backbones, evaluating them using PROCHECK (Laskowski et al., 1993, 1996) and WHAT_CHECK (Hooft et al., 1996) to correct or exclude those not meeting the expectations.
@@
 Side-chains with abnormal torsion angles, atypical aromatic bonding angles, or unusual proline puckering were also discarded.
-Appendix F Molecular Dynamics Simulations
-For experimental structures, crystallographic water and unnecessary small molecules were removed.
+Appendix I Molecular Dynamics and Blind Docking Simulations
+Figure 11: Stability assessment of MD simulations across proteins using various metrics.
+Distributions of (A) RMSD, (B) radius of gyration ($R_{g}$), (C) potential energy, and (D) secondary structure counts throughout the simulation.
+Interquartile ranges and whiskers show metric variation; high-quality structures have medians close to experimentally derived values with narrow ranges.
+For experimentally derived structures, crystallographic water and unnecessary small molecules were removed.
 For generated structures, missing side-chains were added via homology modeling (Section 2.5).
@@
 Simulations were performed with GROMACS (Abraham et al., 2015), using the all-atom CHARMM36 force field (July 2022 version) (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Vanommeslaeghe et al., 2012; Yu et al., 2012; Soteras Gutiérrez et al., 2016).
-F.1 Molecular Dynamics Setup for Stability Assessment of Generated Structures
+I.1 Molecular Dynamics Setup for Stability Assessment of Generated Structures
 […]
@@
 Equilibration involved two stages.
-In the first stage, we performed a 500 ps NVT equilibration (250,000 steps with a 2 fs time step).
+In the first stage, we performed a 500 ps NVT [^6] Constant number of particles, volume, and temperature.
+equilibration (250,000 steps with a 2 fs time step).
 Temperature control was managed using the V-rescale thermostat, with the system divided into two groups: (1) protein and (2) water + ions, both set to a target temperature of 310 K to simulate physiological conditions.
-In the second stage, we carried out a 500 ps NPT equilibration with pressure coupling enabled.
+In the second stage, we carried out a 500 ps NPT  [^7] Constant number of particles, pressure, and temperature.
+equilibration with pressure coupling enabled.
 The pressure was regulated using the C-rescale method with isotropic coupling.
@@
 Full details of the MD parameter files can be found in Software and Data.
-F.2 Molecular Dynamics Setup for Conformational Analysis of Protein-Ligand Complexes
+I.2 Molecular Dynamics Setup for Conformational Analysis of Protein-Ligand Complexes
 The receptor and ligand were saved as separate coordinate files to prepare their respective topologies.
-The receptor topology was prepared as in Section F.1.
+The receptor topology was prepared as in Section I.1.
 For the ligand, hydrogen atoms were added using OpenBabel (O’Boyle et al., 2011), and topology was generated via the CGenFF server (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Vanommeslaeghe et al., 2012).
 The receptor and ligand topologies, along with force-field-compatible coordinate files, were then combined to construct the complete complex system.
-The MD workflow for complexes followed the same steps in Section F.1.
+The MD workflow for complexes followed the same steps in Section I.1.
 Complexes were placed in an octahedral simulation box, energy-minimized in vacuum, solvated in water, and neutralized with Na+ and Cl- ions to 150 mM.
@@
 Additionally, to minimize interference from temperature fluctuations of the ligand on the overall simulation, we defined two temperature coupling groups: (1) the receptor and ligand as one group, and (2) the solvent and ions as the other.
-Other equilibration settings followed Section F.1.
+Other equilibration settings followed Section I.1.
+Figure 12: Comparison of experimentally derived binding modes (bold boxes; first row) with predictions from blind docking simulations for receptors binding to family-specific ligands.
+Ligands are colored as green (ground truth), cyan (predicted using experimentally derived receptors), magenta (predicted using SM samples), and yellow (predicted using FM samples).
+$\Delta G$ (kcal/mol) and RMSD (Å) quantify binding affinities and deviations from the experimental poses.
+“MD” labels protein-ligand complex MD simulations after docking.
 After equilibration, restraints were removed, and a 10 ns production simulation was conducted to analyze the dynamic behavior and conformational changes in the complexes.
-Appendix G Protein-ligand Blind Docking
-Similarly, crystallographic water and unwanted molecules were removed from experimental structures, and missing side-chains were added to generated structures via homology modeling.
+I.3 Protein-ligand Blind Docking
+Similarly, crystallographic water and unwanted molecules were removed from experimentally derived structures, and missing side-chains were added to generated structures via homology modeling.
 Receptor structures were prepared using AutoDock Tools (Morris et al., 2009), with polar hydrogens added, Kollman charges assigned, and any missing atoms repaired.
@@
 The binding mode with the lowest binding free energy was selected as the final result.
-Appendix H Generated Structures
+Appendix J Applicability and Limitations of Deep Generative Protein Design Guidelines
+  • It is hard to apply strict physical constraints to very flexible proteins in generative process.
+In addition, using only steric exclusions without considering environmental factors such as the lipid bilayer in membrane proteins or interfaces in large assemblies can make the designs less accurate (Winnifrith et al., 2024).
+  • If a target protein’s stability or fold depends critically on bound cofactors or oligomeric assembly, designing sequences in isolation can be unreliable.
+According to Krishna et al.
+(2024), generating and screening candidate sequences without accounting for these interactions will often fail to reproduce the native structure or stability observed in the full cofactor- or multimer-associated complex.
+  • For truly new folds or functions, there are no known conserved residues to guide design.
+In families where conserved sites are spread out or have moved over evolution, keeping those sites can block creative changes.
+In practice, the design of novel proteins remains a low-success “attritional” problem with success only in rare, isolated cases (Greener et al., 2018).
+  • Standard force fields often do not model metal ions, sugar attachments, or membrane effects accurately.
+As a result, refinement steps may produce structures that differ from what actually happens in the lab.
+  • For novel protein folds with no known homologs, template-based modeling or comparison is difficult.
+Even state-of-the-art predictors admit substantially reduced accuracy when no homologous structure exists (Jumper et al., 2021).
+  • Conventional MD simulations are limited to relatively short timescales due to computational cost.
+For example, simulating $\sim 50,000$ atoms (a modest protein) for $\sim 1$ $\mu$s can take days on a GPU (Hollingsworth & Dror, 2018).
+Moreover, simulations of systems with membranes, metal centers, or covalent modifications often suffer from force-field artifacts or setup uncertainties.
+  • Docking methods that do not model backbone flexibility often cannot accommodate the large backbone and side-chain rearrangements required for binding, leading to unreliable predictions (Lexa & Carlson, 2012).
+Appendix K Generated Structures
 Figure 13: 50 $\beta$-lactamase-like protein backbones generated using score matching and 50 using flow matching.